### v1

# Multi-modal Retrieval Augmented Multi-modal Generation: A Benchmark, Evaluate Metrics and Strong Baselines

Abstract
This paper investigates an intriguing task of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG). This task requires foundation models to browse multi-modal web pages, with mixed text and images, and generate multi-modal responses for solving user queries, which exhibits better information density and readability. Given the early researching stage of M2RAG task, there is a lack of systematic studies and analysis. To fill this gap, we construct a benchmark for M2RAG task, equipped with a suite of text-modal metrics and multi-modal metrics to analyze the capabilities of existing foundation models. Besides, we also propose several effective methods for foundation models to accomplish this task, based on the comprehensive evaluation results on our benchmark. Extensive experimental results reveal several intriguing phenomena worth further research.
[…]
tcb@breakable
Multi-modal Retrieval Augmented Multi-modal Generation:
A Benchmark, Evaluate Metrics and Strong Baselines
Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
Figure 1: A typical comparison between naive RAG and our proposed M2RAG. GPT-4o is used in this case.
Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the dataset construction pipeline and Step 4 demonstrates the generation process.
1 Introduction
Retrieval Augmented Generation (RAG) has been emerged as a promising technique to enhance the factual correctness of model-generated responses by grounding on facts in external knowledge databases Lan et al. (2023); Sun et al. (2024); Li et al. (2022). Except for textual knowledge, multi-modal data, like images and videos, also contain diverse information, and it is natural to leverage these knowledge to answer user queries more accurately, leading to a concept named Multi-modal RAG or MuRAG Chen et al. (2022); Joshi et al. (2024); Ye et al. (2024). Numerous works Riedler and Langer (2024) have demonstrated that RAG with multi-modal inputs outperforms single-modal RAG. This paper studies MuRAG, but we primarily focus on how to use foundation models to generate multi-modal responses for solving user queries, which we named M2RAG task, given multi-modal inputs with mixed layout of text and images. Numerous previous works Zhu et al. (2024); Wu et al. (2023b); Li et al. (2024) have proven that multi-modal outputs effectively improve the information density and readability for users, embodying the principle—A single image is worth a thousand words. For example, Figure 1 compares plain text and multi-modal responses to the same user query about folding a paper airplane. While both responses provide similar textual explanations, the inclusion of instructional images in the multi-modal response significantly enhances readability. As a result, human readers can understand the instructions more quickly and efficiently, enabling them to successfully create a paper airplane with less effort.
Despite the emerging interest in the M2RAG task, there is a lack of systematic analysis and evaluation, given its early stage of research development. To address this gap, we propose a benchmark designed to evaluate the capabilities of existing generative models, including Large Language Models (LLMs) and Multi-modal Large Language Models (MLLMs), in completing the M2RAG task. This benchmark evaluates the generated content from both text-modal and multi-modal perspectives, incorporating four metrics for each dimension, such as fluency, relevance, image coherence, and helpfulness. The benchmark dataset is constructed through a multi-step workflow, as illustrated in Steps 1-3 of Figure 2: (1) Query Collection: We gather a diverse set of user queries spanning ten distinct topics. For each query, the top-$k$ relevant multi-modal web pages are retrieved using the Google Search API; (2) Data Cleaning: A rigorous data cleaning pipeline is applied to these web pages to extract and filter high-quality text and images, which are then used to build a reliable multi-modal database for retrieval; (3) Retrieval: A suite of retrieval methods is developed to evaluate the relevance between multi-modal data elements and the user query. Each element is assigned a relevance score to guide its selection for further processing. Additionally, we propose two approaches for tackling this task, based on comprehensive studies using the constructed benchmark: (1) Single-stage Approach: This method generates a multi-modal response directly in a single turn, with images represented as placeholders within the response; (2) Multi-stage Approach: This method first generates a text response and then interleaves relevant images, finally refines the text response with retrieved images to ensure a more cohesive and enriched output.
Extensive experimental results uncover numerous intriguing phenomena, providing valuable insights for future researches. For example, the multi-stage approach integrates more relevant images into responses, significantly outperforming the single-stage approach. Moreover, existing MLLMs demonstrate notable limitations in handling M2RAG tasks, performing 8% and 13% worse than LLMs under the single-stage and multi-stage approaches, respectively, which underscores substantial room for improvement. Importantly, compared to separate modeling approaches employed in prior works such as MuRAR Zhu et al. (2024), jointly modeling the relationships and interactions between images and text results in noticeable performance enhancements for the M2RAG task. These findings highlight the necessity of advancing joint modeling techniques to unlock the full potential of multi-modal generative models. Overall, our contributions are summarized as follows:
  • Task Formulation and Benchmark: We introduce the M2RAG task and establish a comprehensive benchmark, alongside a suite of fine-grained text-modal and multi-modal metrics to evaluate the capabilities of existing models.
  • Proposed Baselines: We propose several strong baseline methods for solving the M2RAG task, including single-stage and multi-stage approaches.
  • Experimental Insights: Our experiments demonstrate that current LLMs outperform MLLMs significantly in M2RAG tasks. Additionally, decomposing this complex task into simpler sub-tasks further enhances performance. We also observe that relying solely on images in this task can result in poor performance for certain topics, suggesting that incorporating additional modalities could yield further improvements.
2 Related Work
2.1 Text-modal Retrieval Augmented Generation (RAG)
RAG Li et al. (2022) has been widely used to improve the generation quality of language models by incorporating external knowledge Lan et al. (2023), which is usually retrieved by using BM25, dense retrieval model or search engine Lee et al. (2021); Asai et al. (2023). Recently, given the increasing capabilities of LLMs to effectively process any text inputs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al. (2024).
2.2 Multi-modal RAG
Although typical RAG framework demonstrates great potential, it can only handle textual inputs and cannot utilize the rich information presented in multi-modal data such as images and videos Su et al. (2023). While earlier efforts to incorporate multi-modal inputs for text generation [^1] Most existing multi-modal works mainly focus on the visual modality. often relied on carefully designed frameworks Xiao et al. (2024), a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently Chen et al. (2024); Wang et al. (2024a); Yao et al. (2024); Shen et al. (2023); Riedler and Langer (2024), with representative models including GPT-4o Hurst et al. (2024) and Llama-3.2-Vision Dubey et al. (2024). While these multi-modal models greatly expand the range of applications for LLMs, they remain limited to text-only outputs, restricting their ability to directly deliver rich, multi-modal information to users Zhu et al. (2024).
2.3 Multi-modal Generation
In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts Zhu et al. (2024). Consequently, it is crucial to develop foundation models that not only generate plain text responses to user queries but also incorporate relevant multi-modal data to enhance readability and user engagement. This approach embodies the principle—A single image is worth a thousand words, emphasizing the value of visual elements in effective communication. To the best of our knowledge, MExBERT Singh et al. (2021) took the first step in this research direction, which retrieves one image given the user query and generated response. Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT Wu et al. (2023b), TextBind Li et al. (2024), Janus Wu et al. (2024) and Emu3 Wang et al. (2024b). Unlike these works, our study focuses on M2RAG task, which aims to dynamically select the multi-modal content from multiple multi-modal inputs (text or images) to construct final multi-modal responses without generating any visual elements. To the best of our knowledge, MuRAR Zhu et al. (2024) is the most closely related work to our study. However, it falls short of modeling multi-modal content holistically. Specifically, it generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data. This approach differs significantly from the setup of our proposed M2RAG task, which also emphasizes the integration of multi-modal content in the input.
3 Task Formulation
In M2RAG task, models answer each user query $Q$ with a multi-modal response based on retrieved multi-modal web pages. Suggest that multi-modal data resource $R=\{D_{1},\cdots,D_{n}\}$ represents $n$ retrieved multi-modal documents or web pages from the Internet. Each document $D_{i}$ consists of $m$ ordered elements $\{E_{i,1},E_{i,2},\cdots,E_{i,m}\}$, where each element $E_{i,j}$ can be either a textual or visual element. Elements are preserved in their original order within a document, ensuring that images are accompanied by their corresponding textual contexts. These contexts often provide rich descriptions or additional information that enhances the understanding and relevance of the images.
With the above inputs, the M2RAG task needs several steps to get the multi-modal response. Specifically, the following procedures are essential: (1) In-Doc Retrieval, (2) Generation and (3) Multi-modal Element Incorporation.
In-Doc Retrieval
While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time increases significantly as input sequences grow longer Sun et al. (2024). This challenge is more severe in the M2RAG task, where models must process both textual content and corresponding images simultaneously.
To mitigate this issue, we first propose the In-Doc Retrieval process to collect relevant elements from $R$, which are used for next step multi-modal generation.
As illustrated in Step 3 of Figure 2, we employ a retrieval model, denoted as $M_{R}$, to assess the relevance of each element $E_{i,j}$ in the dataset $R$ with respect to the user query $Q$, denoted as $M_{R}(Q,R)$. The top-$k$ most relevant elements are selected to form a refined multi-modal data resource, $R_{\text{In-Doc}}$. The refined resources serve as the inputs for generative models. Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input data for the generation process.
Generation
The refined multi-modal data resource $R_{\text{In-Doc}}$ is then processed by the generative model $M_{G}$ alongside the user query $Q$ to generate the final multi-modal response:
$r=M_{G}(Q,R_{\text{In-Doc}})$ (1)
Here, $r$ represents the final answer to the user query. Notably, if no visual elements are in $R_{\text{In-Doc}}$, the task defaults to the traditional RAG, since no multi-modal elements are utilized and generated.
Multi-modal Element Incorporation
Different from previous works that retrieve multi-modal data after text generation Zhu et al. (2024), we proposed a novel approach that can incorporate multi-modal data with text in a single-turn conversation. Specifically, images in the input and output are presented as placeholders. Each placeholder follows the following format: ![IMAGE_CONTENT](PSEUDO_URL). The pseudo URL identifies which image the model intends to include Wu et al. (2023a) in the output. For input images, the content is represented as tokenized image embeddings for MLLMs or as detailed image captions for LLMs. For output images, the image content is a concise caption based on the input image content which aids in the Chain-of-Thought (CoT) process Wei et al. (2022), assisting the model in understanding the image content or caption and placing it appropriately.
4 Dataset Construction
As shown in Step 1-3 of Figure 2, we elaborate three detailed steps to construct the benchmark dataset for evaluating the capabilities of foundation models in M2RAG task: (1) Query Collection, (2) Data Preparation and (3) Element Evaluation.
4.1 Query Collection
The queries are sourced from the ELI5 Fan et al. (2019) dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2. Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for providing comprehensive answers. After filtering, the remaining queries are classified to ensure a balanced dataset and lay the groundwork for further studies on the topic sensitivity of the proposed methods [^2] Please refer to Appendix A for more details.. The classification process results in 11 distinct categories, as detailed in Table 1. 20 queries are sampled in each category (other than “Others”), resulting in a final benchmark dataset containing 200 queries.
ID Category Name Num. Samples
1 Society & Culture 14182
2 Science & Mathematics 14449
3 Health 10931
4 Education & Reference 776
5 Computers & Internet 6999
6 Sports 1974
7 Business & Finance 3821
8 Entertainment & Music 1946
9 Family & Relationships 820
10 Politics & Government 4638
11 Others 451
Table 1: Topics and volumes of the queries.
4.2 Data Preparation
All multi-modal data used for retrieval and generation are prepared during this step. Specifically, data for each query are processed through the following procedures: (1) Prepare Web Pages and Auxiliary Images, (2) Text Processing and (3) Image Processing.
4.2.1 Prepare Web Pages and Auxiliary Images
For each collected query, we first employ Google Search API [^3] https://developers.google.com/custom-search/v1 to get the URLs of web pages given the query, and then retrieve each web page containing a mix of text and images in markdown format using JINA AI Reader [^4] https://jina.ai/reader/. Our preliminary study reveals that the images within these web pages are sometimes inaccessible, low quality, or insufficient for effective multi-modal generation. Therefore, we supplement the dataset with auxiliary images sourced from Google Image Search, using the user queries as search input to ensure relevance and enrich the visual content.
Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset, as illustrated in Step 2 of Figure 2.
4.2.2 Text Processing
The raw web pages often contain a significant amount of irrelevant or redundant text, like the advertisements and web links, and it is essential to remove these data and produce concise and relevant content for generation. The text processing process is divided in to three steps: (1) Embedded Image Handling: image URLs are extracted from the web pages for image collection, and placeholders are inserted in their place within the text; (2) Text Cleaning: text snippets matching pre-defined patterns, such as web links, are removed to ensure cleaner and more relevant content; (3) Text Splitting: the cleaned text content is segmented into smaller pieces using the LangChain Chase (2022) text splitter, facilitating further selection and processing.
4.2.3 Image Processing
We conduct following steps to process images, ensuring they are of high quality and relevant to the user query: (1) Caching & Conversion: we download all the images with their URLs, and convert them into widely accepted formats such as JPG, PNG, GIF, or WEBP. Images that fail to be downloaded or converted are discarded; (2) Filtering: images that fall below a certain size threshold or receive a low similarity score with CLIP Radford et al. (2021) are removed. Such images are typically insufficiently informative, often consisting of icons, banners, or other non-representative visuals. (3) Deduplication: PHash algorithm-based deduplication Zauner (2010) is applied to remove duplicated or similar images, preventing redundant images for generation.
4.3 Element Evaluation
As described in Section 3, we need to perform In-Doc retrieval to select text and visual elements for efficient generation. To achieve this goal, the relevance between each element and the user query should be evaluated. Specifically, we utilized the DeepSeek-Chat DeepSeek-AI (2024) API for evaluating textual elements and MiniCPM-V-2.6 Yao et al. (2024) for assessing visual elements. Our preliminary manual annotations revealed that these methods demonstrate a high degree of consistency with human evaluations in measuring correlation, making them suitable for this task [^5] Please refer to Appendix B for more details.. This step results in a score from 0 to 10 for each element, where higher scores indicate better relevancy. This can be used to replace the traditional embedding-based similarity score for retrieval.
With the above procedures, a high-quality multi-modal RAG dataset with real-world user queries, retrieved web pages and images is constructed. Detailed statistical information of our constructed dataset can be found in Appendix C.
5 Strong Baselines for M2RAG Task
Finishing Step 1-3 of Figure 2, we introduce three strong baselines for Step 4 to accomplish our proposed M2RAG task.
5.1 Separate Modeling
Previous studies have approached text generation and image integration as distinct processes, exemplified by the MuRAR method Zhu et al. (2024). MuRAR operates in three sequential steps: (1) retrieving relevant plain text content and generating a text-based response; (2) identifying and incorporating appropriate visual elements into the text-based response; and (3) refining the text to seamlessly integrate the visual elements and their associated information.
In our implementation, pre-trained embedding model Reimers and Gurevych (2019) is used to retrieve visual elements based on their surrounding text content or detailed image caption generated by GPT-4o. Since the separated method does not need to process multi-modal content, it is implemented by LLMs.
5.2 Joint Modeling
The separate approach overlooks the relationships between textual content and multi-modal data when generating text responses, leading to sub-optimal integration of multi-modal information. To address this limitation, we propose several baseline methods that jointly model text and images as a unified entity, capturing their rich relationships and interactive dynamics.
Specifically, our approach can be categorized into single-stage and multi-stage methods, leveraging LLMs or MLLMs as backbone models. This design allows for a more holistic representation of multi-modal content, as depicted in Step 4 of Figure 2.
Single- and Multi-stage Approaches
Single-stage approach requires models to directly generated a multi-modal output with selected images in their corresponding placeholders, given all multi-modal content saved in a single prompt.
Although the single-stage approach is straightforward, current open-source foundation models struggle in generating accurate image placeholders. Besides, it is a well-known limitation that most MLLMs may struggle with handling a large number of images simultaneously. Therefore, we propose a multi-stage approach to address the complexity of the single-stage approach, which involves three stages: (1) Text Response Generation: generate a plain text answer given the multi-modal input; (2) Image Interleaving: divide the text into several segments, and LLMs or MLLMs are prompted to determine which segment requires the inserted images for better readability; (3) Text Refinement: refine each segment by incorporating the information of selected images. Although multi-stage approach addresses the shortcomings of single-stage approach, their computational overhead for inference is much higher.
LLM and MLLM Backbones
Both of LLMs and MLLMs can be used to solve M2RAG task with single-stage or multi-stage approaches. Specifically, as introduced in Section 3, given the retrieved multi-modal data resource $R_{\text{In-Doc}}$, the prompt $P$ for LLMs and MLLMs could be constructed using a structured template $T$:
$P=T(G,Q,R_{\text{In-Doc}})$ (2)
where $G,Q$ denotes the textual task guidelines and user query, respectively. Then LLMs and MLLMs could be prompted with $P$ to solve M2RAG task: (1) For MLLMs, the encoded image tokens are placed at their original positions  Dubey et al. (2024); Wang et al. (2024a), reserving the relative position in the multi-modal web pages. Since the additional images do not belong to any web pages, their image tokens are placed in a distinct section in the prompt; (2) For LLMs, since they could not process images directly, the detailed captions of images in $R_{\text{In-Doc}}$ are provided, allowing LLMs to access the semantical content of images.
6 Evaluation Metrics
To evaluate the effectiveness of numerous models and methods, we design four text-modal metrics and four multi-modal metrics. Since the generated response are open-ended, these evaluation metrics are evaluated by prompting advanced GPT-4o model in a reference-free manner Lan et al. (2024). To the best of our knowledge, prompting GPT-4o for automatic evaluation is the most reliable automatic evaluation methods for open-ended RAG and multi-modal evaluation scenarios Es et al. (2023); Ku et al. (2024); Lu et al. (2023). The detailed prompts for our proposed text-only and multi-modal evaluation metrics can be found in Appendix D. The score ranges of all fine-grained text-only and multi-modal metrics are from 0 to 1, and scores close to 1 indicates better performance.
6.1 Text-modal Metrics
Following previous works Zhu et al. (2024), we utilize both linguistic and RAG metrics Es et al. (2023) to assess the quality of the generated content in the M2RAG task. These metrics ensure a comprehensive evaluation of the generated textual content.
Fluency
assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al. (2022); Es et al. (2023).
Relevance
assesses how pertinent the model-generated textual content is to the given user query Lan et al. (2022); Es et al. (2023).
Context Precision
evaluates the proportion of relevant chunks in the retrieved web pages by analyzing the overlap of their key concepts Es et al. (2023).
Faithfulness
measures the accuracy of the model-generated content in representing information from the original materials, focusing on factual alignment to avoid fabricated or misleading details Es et al. (2023); Sun et al. (2024).
6.2 Multi-modal Metrics
In addition to text-only metrics, we also focus on the quality of model-generated multi-modal content from following five perspectives.
Image Coherence
examines the logical and coherent alignment of images with their surrounding text content, ensuring visual content enhances and complements the textual narrative Zhu et al. (2024).
Image Helpfulness
evaluates how effectively images contribute to the user’s comprehension of the text. This involves assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al. (2024).
Image Reference
verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading.
Image Recall
measures whether the generated content includes highly relevant, informative, and important images. As described in Section 4.3, images are evaluated and scored during pre-processing, with selection based on these scores prior to generation. Consequently, the images input to generative models are generally relevant to the user query. However, issues such as semantic overlap or quality variations among images may arise, leading to the inclusion of sub-optimal images in the retrieved images. The goal of Image Recall is to ensure that no images other than the sub-optimal ones are excluded from the final answer. To achieve this, we prompt MLLMs to assess whether each image enhances human understanding of the content. To address potential semantic overlap, we include textual descriptions of images already present in the document and instruct the model to classify an image as negative if its content is significantly redundant compared to existing images. This process ensures that only diverse and meaningful images are retained.
6.3 Overall Score Computation
Except for these fine-grained evaluation dimensions, it is essential to summarize a final overall score reflecting the capabilities of foundation models in M2RAG task. Specifically, the average of four text-only metrics and five multi-modal metrics is computed as the overall score.
7 Experimental Results
Model Type Approach Model Text-modal Metrics Multi-modal Metrics Overall
Flu. Rel. CP. Faith. Coher. Help. Ref. Recall
LLMs Separate GPT-4o 0.80 0.82 0.69 0.89 0.53 0.44 0.43 0.43 0.63
Single GPT-4o 0.79 0.81 0.76 0.91 0.69 0.57 0.31 0.57 0.68
Llama-3.1-70B-Instruct 0.75 0.79 0.75 0.85 0.57 0.47 0.18 0.22 0.57
Qwen2.5-72B-Instruct 0.79 0.83 0.77 0.87 0.62 0.52 0.35 0.52 0.66
Llama-3.1-8B-Instruct 0.74 0.77 0.72 0.83 0.49 0.38 0.28 0.62 0.60
Qwen2.5-7B-Instruct 0.71 0.73 0.73 0.91 0.52 0.41 0.22 0.63 0.61
Average 0.75 0.79 0.75 0.87 0.56 0.45 0.27 0.51 0.62
Multi GPT-4o 0.78 0.76 0.68 0.87 0.68 0.60 0.81 0.97 0.77
Llama-3.1-70B-Instruct 0.73 0.75 0.72 0.84 0.68 0.55 0.77 0.97 0.75
Qwen2.5-72B-Instruct 0.77 0.77 0.72 0.85 0.69 0.58 0.78 0.97 0.77
Llama-3.1-8B-Instruct 0.72 0.73 0.72 0.82 0.66 0.55 0.75 0.93 0.74
Qwen2.5-7B-Instruct 0.74 0.74 0.74 0.83 0.64 0.55 0.76 0.95 0.74
Average 0.75 0.75 0.72 0.84 0.67 0.57 0.77 0.96 0.75
MLLMs Single GPT-4o 0.78 0.82 0.76 0.90 0.64 0.53 0.24 0.60 0.66
Llama-3.2-90B-V-Instruct 0.77 0.68 0.64 0.76 0.51 0.34 0.10 0.01 0.48
Qwen2-VL-72B-Instruct 0.77 0.78 0.72 0.88 0.43 0.32 0.15 0.16 0.53
Llama-3.2-11B-V-Instruct 0.77 0.64 0.60 0.68 0.26 0.21 0.02 0.02 0.40
Qwen2-VL-7B-Instruct 0.66 0.64 0.71 0.83 0.39 0.30 0.23 0.15 0.49
Average 0.75 0.71 0.69 0.81 0.53 0.43 0.22 0.19 0.54
Multi GPT-4o 0.77 0.78 0.72 0.86 0.65 0.56 0.81 0.97 0.76
Llama-3.2-90B-V-Instruct 0.72 0.70 0.67 0.74 0.55 0.43 0.62 0.94 0.67
Qwen2-VL-72B-Instruct 0.75 0.74 0.71 0.80 0.63 0.52 0.73 0.94 0.73
Llama-3.2-11B-V-Instruct 0.72 0.69 0.70 0.69 0.47 0.34 0.34 0.81 0.59
Qwen2-VL-7B-Instruct 0.70 0.75 0.73 0.77 0.54 0.42 0.51 0.94 0.67
Average 0.73 0.73 0.70 0.77 0.57 0.46 0.62 0.92 0.69
Table 2: The overall experiment results on M2RAG task. Flu., Rel., CP. and Faith. represent fluency, relevance, context precision and faithfulness, respectively. Coher., Help., Ref., and Recall represent image coherence, helpfulness, reference and recall, respectively. The highest scores for each group are highlighted in bold, and the highest average scores for 4 groups are highlighted in red.
Using the proposed baseline methods and evaluation metrics, we conduct comprehensive experiments on the benchmark dataset with a range of representative models, including GPT-4o, the Llama-3.1 series, Qwen2.5 series, Llama-3.2-Vision series, and Qwen2-VL series [^6] Please refer to Appendix E for more details.. In this section, we analyze the experimental results presented in Table 2 from the following perspectives.
The Most Capable Models
As shown in Table 2, it can be found that the most capable open-source and closed-source foundation models for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o (0.77), respectively, under the experimental setup of multi-stage approach with LLM backbones. Although the performance gap between these two models are negligible with single-stage approach, there is still a noticeable gap when it comes to single-stage approach(0.68 > 0.66). For MLLM backbones, GPT-4o significantly outperforms all closed-source models. For example, Qwen2-VL-72B-Instruct is much worse than GPT-4o by using the single-stage approach (0.53 < 0.66). Besides, there is still significant room for improvement in the performance of the most capable GPT-4o model (0.77 < 1.0) in the M2RAG task.
Scaling Phenomena
Significant scaling phenomena are observed in most experimental setups, with model performance improving as model size increases. However, a notable exception is found in the single-stage approach with LLMs. For instance, Llama-3.1-8B-Instruct slightly outperforms Llama-3.1-70B-Instruct (0.60 > 0.57) under this approach. Rigorous human annotation reveals that this issue arises because larger models tend to select fewer images in the single-stage approach. It should be noted that many of the excluded images are beneficial for the response, and their omission results in performance degradation. For example, the Image Recall scores of Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct are significantly lower than those of Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct (0.22 < 0.62 and 0.52 < 0.63, respectively). This phenomenon highlights a limitation in the image selection behavior of larger models, which warrants further investigation in future work.
Joint Modeling versus Separate Modeling
It can be observed that the separate modeling method is much worse than all of our proposed approaches. Specifically, our worst approach, i.e., single-stage with MLLMs (GPT-4o) still outperforms it (0.66 > 0.63). This phenomenon demonstrates that jointly modeling the relationship and interactive information between multi-modal data is crucial for the M2RAG task.
Comparison among Approaches
As shown in the “Average” rows in Table 2, it can be found that the multi-stage approach significantly outperforms the single-stage approaches for LLMs models (0.75 > 0.62) and MLLMs models (0.69 > 0.54). Besides, multi-stage approach introduce more relevant images, enhancing the information density and readability of the generated answers. Specifically, the Image Recall and Helpfulness fine-grained scores of multi-stage approaches are higher than the single-stage approach.
Moreover, the LLMs significantly outperforms the MLLMs in M2RAG tasks, no matter the single-stage or multi-stage is used. For example, the average overall score of LLMs are better than MLLMs (0.71 > 0.64). This observation demonstrates that the capabilities of existing MLLMs are not yet capable enough to accomplish the complex M2RAG tasks.
Model Approach Prompt Completion Total
LLMs Single 7797 767 8563
Multi 16947 2252 19199
MLLMs Single 11454 553 12007
Multi 24078 1707 25785
Table 3: Average token usage for different approaches.
Furthermore, we also compute the average cost, i.e., the average prompt and completion tokens of GPT-4o, in Table 3, and we could summarize two conclusions: (1) the cost multi-stage approach is over 2 times that of single-stage; (2) the cost of MLLMs is approximately 1.3-1.4 times that of LLMs. Therefore, considering that LLMs generally performs better than MLLMs in M2RAG, using single-stage with LLMs as backbone model is currently the most cost-effective approach.
8 Analysis
In this section, we answer the following three important questions about our proposed approaches.
8.1 Topic is a Key Factor Affecting Performance
The average overall scores of all evaluated models, as presented in Figure 3, reveal notable variations in performance across different topics. For instance, the most challenging topic is “Business & Finance”, which achieves the lowest overall score, whereas “Society & Culture” and “Sports” are significantly easier topics by comparison.
Figure 3: Average overall score across 10 topics.
We hypothesize that these performance differences stem from the varying demands for multi-modal information across topics. Specifically, images are more prevalent and useful in less technical fields, such as sports and culture. In contrast, fields like business and finance often require data from additional modalities, such as tables and charts, which are less readily integrated. Moreover, theoretical fields like politics may not rely heavily on multi-modal information for comprehensive responses.
To achieve balanced performance across topics, a topic-sensitive approach should be developed to adaptively incorporate appropriate modalities for response synthesis, ensuring that the specific requirements of each topic are effectively addressed.
8.2 Caption Quality is Important for Image Integration
As described in Section 5.2, images are converted into captions for use by LLMs. To ensure a comparable level of information between LLMs and MLLMs, the captions are crafted to be as detailed as possible. Additionally, textual contexts associated with the images are included during the captioning process to enhance consistency. In this section, we investigate the impact of removing these elements.
Figure 4: Comparison of performance with original and ablation settings.
As shown in Figure 4, performance declines noticeably with the removal of details and contextual information, particularly in the metrics of Image Coherence and Image Reference. The absence of contextual information in captions leads to a significant drop in Image Recall, highlighting the critical role of context in guiding the image selection process of LLMs. Beyond these key observations, the ablation settings show minor fluctuations in other metrics. Overall, incorporating context during captioning is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
8.3 Contributions of Auxiliary Images
Auxiliary images are collected for each query to address cases where highly relevant images are absent from the retrieved web pages. Removing these images leads to a significant decline in generative performance (Figure 4). For instance, Image Recall drops sharply without auxiliary images, suggesting that LLMs preferentially include auxiliary images over those sourced from web pages when constructing the generated document. Additionally, the decline in Image Coherence and Image Helpfulness metrics highlights the superior quality of auxiliary images compared to web-sourced images. Interestingly, the Image Reference metric shows a notable increase, likely due to stronger connections between each remaining image and its context. Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
9 Conclusion
In this paper, we formate a challenging task—M2RAG, requiring foundation models to process the mixed multi-modal web pages to generate a multi-modal answers for solving user queries. Besides, we construct a benchmark to comprehensively evaluate the capabilities of existing foundation models based on four text-only and five multi-modal fine-grained metrics. Furthermore, we also propose several strong baselines for existing models to solve this task. Extensive experimental results demonstrate several intriguing phenomena, facilitating the future research in this field.
10 Limitations
Although our proposed methods outperform some existing approaches for the M2RAG task, there are still several limitations.
Topic Restrictions
Certain topics, such as “Business & Finance” or “Politics & Government”, are not well-suited for visual representation. For these domains, tables, charts, or other formats may be more appropriate than images. The absence of these modalities limits the applicability of our approach to these areas.
Content Availability
The lack of extensive public discussion or data in certain specialized fields presents challenges for retrieving abundant multi-modal documents effectively. This limitation reduces the comprehensiveness of our analysis in underrepresented domains.
Evaluation Metrics
The proposed evaluation metrics are limited and may fail to capture all aspects of M2RAG evaluation. For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer. However, using the LLM-as-a-Judge Zheng et al. (2023) approach, pairwise comparisons are resource-intensive, particularly for large-scale evaluations. Furthermore, their high cost and impracticality make them unsuitable for benchmarking, necessitating reliance on fixed scoring metrics that may not adequately address all edge cases.
Appendix A Query Collection
To simulate the real-world user query solving problem, we collect diverse and high-quality user queries from the ELI5 dataset Fan et al. (2019). The ELI5 corpus is particularly suited for our task due to its comprehensive collection of long-form, open-ended questions that necessitate detailed, multi-sentence responses. The diversity of topics in ELI5 presents an opportunity to reflect the capabilities of language models in the real-world scenarios. Subsequently, we conduct two steps to collect diverse and high-quality queries that are suitable for using multi-modal information as responses: (1) Query Filtering; and (2) Query Classification.
# Task Description
You are a useful text classification assistant. You will be provided with a piece of text to be recognized. Your task is to classify the text with the guidelines.
# Guidelines
1. If the input text is not a question, the description is unclear, you are supposed to generate a single integer 0, which means it is an invalid question.
2. If there is no need to cite rich references to answer the question, i.e. the question can be fully addressed with a simple answer, you are supposed to generate a single integer 0, which means it is an invalid question. 3. If the question is not classified as invalid question according to the above two rules, then generate a single integer 1. # Output Format
Your output should consist of two lines.
- The first line is your brief analysis of the input question content and why or why not it meets the requirements; - The second line is an single integer 0 or 1. Do not generate any other information other than the analysis and category indices. # Input Text
{text}
# Output
Figure 5: Prompt template for filtering queries which are not complex questions.
# Task Description
You are a useful question classification assistant. Your task is to classify the input question with the guidelines.
# Guidelines
1. If inserting appropriate images in the answer to this question can provide a more comprehensive answer and make it easier for users to understand the content being discussed, then generate a single integer 1.
2. Otherwise, generate a single integer 0. # Output Format
Your output should consist of two lines.
- The first line is your brief analysis of the input question content and why or why not it meets the requirements; - The second line is an single integer 0 or 1. Do not generate any other information other than the analysis and category indices. # Input Question
{question}
# Output
Figure 6: Prompt template for filtering queries which are not necessarily answered with images.
Query Filtering
conduct two steps to filter user queries that require detailed explanation and visual information for better understanding the content in the responses. The first step involves eliminating invalid or low-quality queries. Specifically, GPT-4o is prompted to remove queries that do not elicit detailed responses, such as those reducible to yes/no answers or simple sentences (as illustrated in Figure 5). This ensures that the dataset focuses on queries requiring more elaborate, informative responses. The second step filters out queries that can be fully addressed through textual content alone, as illustrated in Figure 6. Queries that lack the necessity for visual information do not align with the goal of evaluating multi-modal capabilities and are thus excluded.
# Task Description
You are a useful question classification assistant. Your task is to classify the input question with the guidelines.
# Category List
1. Society & Culture
: This category encompasses questions about traditions, languages, customs, behaviors, and societal norms. Images can serve as a visual reference to cultural landmarks, traditional attire, or historical events that support the textual explanation.
Example: What is the significance of the Day of the Dead celebration in Mexican culture? 2. Science & Mathematics: Questions in this category often deal with complex concepts that can be enhanced with diagrams, charts, or images. These visual aids can help in understanding scientific phenomena or mathematical theories. Example: How does photosynthesis work in plants? 3. Health: Health-related questions often benefit from the use of images like anatomical diagrams, charts, or photos for better understanding of medical conditions, treatments, or fitness exercises. Example: What are the symptoms and treatment options for carpal tunnel syndrome? 4. Education & Reference: This category includes academic topics and general knowledge where explanatory images, instructional graphics, and reference tables can provide clarity and enhance learning. Example: What are the main parts of a plant cell and their functions? 5. Computers & Internet: Questions about technology, software, and internet usage can be more effectively answered with screenshots, infographics, and step-by-step visual guides. Example: How do I set up a Wi-Fi network at home? 6. Sports: Sports-related queries can be explained better with visual demonstrations, play diagrams, athlete photos, and equipment images to illustrate techniques, rules, or historical moments. Example: What are the basic rules of soccer? 7. Business & Finance: Financial and business concepts often involve data, charts, and visualizations that can make complex information more digestible and insightful. Example: How does compound interest work and what impact can it have on investments? 8. Entertainment & Music: This category includes questions about movies, TV shows, music, and celebrities where images of album covers, scene snapshots, or charts of music theory can enrich the content. Example: What was the cultural impact of the TV show ’Friends’?
9. Family & Relationships: Questions about family dynamics, parenting, and social relationships often benefit from illustrative images such as family diagrams, photos, and infographics about communication techniques or social scenarios. Example: What are some effective conflict resolution strategies for couples? 10. Politics & Government: Political and governmental questions can be clarified with the help of charts, maps, and historical photographs that provide visual context to legislative processes, electoral maps, or political events. Example: How does the electoral college system work in the United States? 11. Others
# Classification Guidelines
Select one or more categories from categories 1-10 that are most relevant to the content of the question as the category of the question. If the question does not belong to any of the categories 1-10, then classify it into category 11.
# Precautions
1. If the input text is classified into category 11, it cannot be assigned any other category.
2. Please first briefly analyze the content of the input text and find its association with the given category, and then output category indices based on the above analysis. # Output Format
Your output should consist of two lines.
- The first line is your brief analysis of the input text content and its relationship with the given category; - The second line consists of a list of integers separated by a space. Each integer represents the index of a category you assigned to the input question. Do not generate any other information other than the analysis and category indices. # Input Question
{question}
# Output
Figure 7: Prompt template for query classification.
Query Classification
To effectively analyze the difference among user query topics, we also prompt GPT-4o model to automatically categorize these queries into eleven topics from Yahoo Answers Topics [^7] https://huggingface.co/datasets/community-datasets/yahoo_answers_topics. Table 1 demonstrates that our user queries covers common and diverse searching scenarios, which is beneficial to reflecting the performance of evaluated models in real-world scenarios. The prompt template for query classification is illustrated in Figure 7.
Appendix B Retrieval of Collected Elements
We replaced traditional embedding-based retrieval methods with LLM-based and MLLM-based approaches to enhance the retrieval process. Embedding-based methods typically prioritize semantic similarity but often overlook whether candidate elements provide substantial information relevant to the target question. Additionally, in real-time M2RAG tasks, where online documents cannot be pre-indexed, LLM-based and MLLM-based methods offer better performance without incurring significant additional time costs.
Using these methods, text and visual elements within the collected documents are evaluated based on their relevance to the user query. This evaluation is conducted by prompting LLMs and MLLMs for textual and visual content, respectively. The prompt templates used for these evaluations are illustrated in Figures 8 and 9.
# Task Description
You are a useful text evaluation assistant. You will be provided with a user query and a piece of text extracted from a relevant web page. Please evaluate whether this text contains useful information to address the user query.
# Input Data
1. User Query
2. Text from a Web Page # Evaluation Guidelines
Assign a score from 0 to 10, where a higher score indicates better alignment:
- 0: Entirely irrelevant to the query. - 1-3: Minimal relevance; most content deviates from the query. - 4-6: Moderately relevant; some content aligns with the query, but includes notable digressions. - 7-9: Highly relevant; most content relates directly to the query with minor digressions. - 10: Perfectly relevant; fully aligns with the query without any digressions. # User Query
{query}
# Web Page Text
markdown
{webpage_piece} ‘‘‘ # Precautions
If there is an <IMAGE_PLACEHOLDER> in the webpage text, ignore it in your evaluation.
# Output Format
Your output should consist of two lines:
1. A brief analysis of the text’s relevance to the query. 2. An integer score from 0 to 10. # Output
Figure 8: Prompt template for the retrieval of text elements.
# Task Description
You are an assistant for evaluating the correlation between image and text. You will be provided with an image and a question. The image comes from a web page, so it may be a highly relevant image to the provided text, a slightly relevant image, or a completely unrelated image, or even an advertising image.
# Scoring Strategy
Score from 0 to 10 based on the correlation between the image content and the problem, with higher scores indicating higher correlation.
- 0 represents that the image is completely unrelated to the question, or the image is an advertising image. - 1-3 represents that the image is slightly related to the question, and is dispensable for answering the question. - 4-6 represents that the image is moderately related to the question, and it is helpful for answering some of the content in the question. - 7-9 represents that the image is highly related to the question, and it is very helpful for answering the question. - 10 represents that the image is the key to the question, and it is essential for answering the question. # Precautions
You are not expected to generate any other information other than the score (an integer).
# Question
{question}
# Score
Figure 9: Prompt template for the retrieval of visual elements.
Appendix C Dataset Statistics
C.1 Basic Statistics
The statistical information about our benchmark are shown in Table 4.
Item Value
Queries 200
Web pages per query 6.4
Text elements per web page 7.2
Images per query 9.3
Auxiliary image per query 2.6
Average score of text elements 1.1
Average score of image elements 4.2
Average auxiliary image score 3.6
Table 4: Statistical information of our benchmark.
C.2 Data Distribution
We list the distributions of some key items:
  • The number of non-empty web pages for each query, as illustrated in Figure 16;
  • The number of valid text pieces in each web page, as illustrated in Figure 16;
  • The score of each text piece, as illustrated in Figure 16;
  • The number of images in web pages for each query, as illustrated in Figure 16;
  • The score of each web page image, as illustrated in Figure 16;
  • The number of auxiliary images for each query, as illustrated in Figure 16;
  • The score of each auxiliary image, as illustrated in Figure 16.
Figure 10: Non-empty web pages for each query
Figure 11: Valid pieces in each web page
Figure 12: Scores of web page pieces
Figure 13: Web page images for each query
Figure 14: Score of web page images
Figure 15: Auxiliary images for each query
Figure 16: Score of auxiliary images
Appendix D Details of Evaluation Metrics
Apart from Context Precision and Faithfulness which are implemented with RAGAS, other metrics are implemented by prompting GPT-4o with the following prompt templates:
  • Fluency: as illustrated in Figure 17;
  • Relevance: as illustrated in Figure 18;
  • Image Coherence: as illustrated in Figure 19;
  • Image Helpfulness: as illustrated in Figure 20;
  • Image Reference: as illustrated in Figure 21;
  • Image Recall: as illustrated in Figure 22.
# Task Description
You are a document evaluation assistant. Your task is to evaluate the fluency of the provided text written in markdown.
# Input Text
markdown
{text} ‘‘‘ # Scoring Criteria
Assign a score from 0 to 10 based on the fluency of the text:
- 0: The text is incoherent, lacks logical flow, and deviates significantly from correct markdown syntax. - 1-3: The text has potential with partial fluency and logic but is plagued by multiple markdown errors, such as unhandled duplicate headings or incorrect formatting structures. - 4-6: The text demonstrates general fluency with minor grammatical or logical inconsistencies, though the markdown structure lacks clarity or uniformity, like having redundant sections. - 7-9: The text is well-written and logically coherent, with proper markdown usage, forming a mostly seamless document that maintains consistent quality. - 10: The text is exemplary, demonstrating perfect fluency, logical progression, and impeccable markdown syntax, representing an ideal markdown document. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
1. A brief explanation justifying the assigned score. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 17: Prompt template for the text fluency evaluation.
# Task Description
You are a document evaluation assistant. Your task is to evaluate the relevance between the query and provided text written in markdown.
# Query
{query}
# Input Text
markdown
{text} ‘‘‘ # Scoring Criteria
Assign a score from 0 to 10 based on the degree of relevance between the input text and the given query:
- 0: The text is entirely irrelevant to the given query. - 1-3: The text has minimal relevance, with most content deviating from the intended query. - 4-6: The text is moderately relevant, with some content aligned with the intended query, but contains noticeable digressions. - 7-9: The text is highly relevant, with most content directly related to the intended query and only minor digressions. - 10: The text is perfectly relevant, fully aligning with the intended query without any digressions. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
1. A brief explanation justifying the assigned score. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 18: Prompt template for the text relevance evaluation.
# Task Description
You are a multi-modal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the coherence between the image and the text (context above and below) it accompanies.
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria
Assess how coherent the image is in relation to its accompanying text, assigning a score from 0 to 10. A higher score indicates stronger coherence between the image and the text. Be precise when assigning the score.
- A score from 0-3 means that the image is minimally or not at all coherent with the text. - A score from 4-6 indicates that the image shows some coherence with the text but may include unrelated elements. - A score from 7-9 indicates that the image is highly coherent with the text. - A score of 10 indicates perfect coherence, where the image completely corresponds with and enhances the text. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
1. A brief explanation justifying the assigned score. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 19: Prompt template for the image coherence evaluation.
# Task Description
You are a multi-modal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the helpfulness of the image in enabling human readers to comprehend the text (context above and below) it accompanies.
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria
Evaluate how well the image helps human readers understand the content of its accompanying text, assigning a score from 0 to 10. A higher score indicates that the image significantly enhances comprehension of the text. Be as precise as possible when assigning the score.
- 0-3: The image is minimally or not at all helpful for comprehension. - 4-6: The image provides some helpful context or information but may contain extraneous or less relevant details. - 7-9: The image is highly helpful in enabling comprehension of the text. - 10: The image is perfectly helpful and greatly enhances and clarifies the information provided in the text. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
1. A brief explanation justifying the assigned score. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 20: Prompt template for the image helpfulness evaluation.
# Task Description
You are a multi-modal document quality assessment assistant. You will receive an image and its accompanying textual context, formatted in markdown. Your task is to determine whether the image is explicitly referenced or explained within the surrounding text (both above and below the image).
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria
Evaluate the extent to which the image is referenced or explained in the text:
- 0: The image is not mentioned or referenced in the context. - 1-3: The image is referenced implicitly, and the reference is improper or incorrect. - 4-6: The image is referenced explicitly but in an improper manner, or it is referenced implicitly. - 7-9: The image is referenced explicitly, with the reference being generally proper and correct. - 10: The image is referenced explicitly, with the placement and explanation being completely proper and correct. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should consist of only two lines:
1. A brief explanation justifying the assigned score. 2. The score as an integer value. Refrain from providing any additional information beyond the explanation and score.
Figure 21: Prompt template for the image reference evaluation.
# Task Description
You are a multi-modal document evaluation assistant. You will receive a single image and a multi-modal document written in markdown, where detailed captions serve as proxies for images. Your task is to evaluate whether the image should be inserted into the document to enhance understanding for human readers.
# Images in the Document
The images in the document have been replaced with their detailed captions. Each image is represented by a single line starting with ‘<image>:’.
# Document
markdown
{text} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Scoring Criteria
Evaluate whether the provided image should be inserted into the document based on the following:
- If the image is irrelevant to the document’s content or does not aid in comprehension, output a single integer 0, indicating the image should not be inserted. - If the image content is too similar to or repetitive of images already represented in the document, output a single integer 0, indicating the image should not be inserted. - Otherwise, output a single integer 1, indicating the image should be inserted. # Output Instructions
Your output should contain only two lines:
1. A brief explanation justifying the score of 0 or 1. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 22: Prompt template for the image recall evaluation.
Appendix E Experimental Setup
Evaluated Models
We evaluate the capabilities of some advanced open-source and closed-source LLMs and MLLMs for M2RAG task. The involved models are listed in Table 5. Besides, we also implement a separate approach, i.e., the MuRAR method with GPT-4o as the backbone model Zhu et al. (2024).
LLM MLLM
GPT-4o GPT-4o
Llama-3.1-70B-Instruct Llama-3.2-90B-Vision-Instruct
Llama-3.1-8B-Instruct Llama-3.2-11B-Vision-Instruct
Qwen-2.5-72B-Instruct Qwen-2-VL-72B-Instruct
Qwen-2.5-7B-Instruct Qwen-2-VL-7B-Instruct
Table 5: LLMs and MLLMs used for M2RAG task.
Implementation Details
For the generation process, we deployed open-source LLMs and MLLMs using vLLM Kwon et al. (2023) on Nvidia A100-SXM4-80GB GPUs. Models with 7B-11B parameters operate on a single GPU, while those exceeding 70B parameters are distributed across 4 GPUs using tensor parallelism. The context length is configured at 64k for the Llama series and 32k for the Qwen series. During generation, we use a top-$k$ selection method with $k$ set to 20 for filtering text elements from each webpage. The maximum number of auxiliary and web page specific images is restricted to 5, with a total input image cap at 10. Images are resized to 512 $\times$ 512 thumbnails for MLLMs processing. For evaluation, we implemented our customized metrics using the GPT-4o model to ensure the robust comprehensive assessment. Besides, the RAG metrics are implemented with RAGAS Es et al. (2023) by using the GPT-4o-mini model.
Appendix F Case Study
In this section, we provide some cases to demonstrate more details in the M2RAG task:
  • Single-stage versus Multi-stage, as illustrated in Figure 23;
  • Large Models versus Small Models, as illustrated in Figure 24;
  • Separate Modeling versus Joint Modeling, as illustrated in Figure 25.
Figure 23: A case of comparison between single- and multi-stage approaches with Llama-3.2-Vision-90B-Instruct.
Figure 24: A case of comparison between small and large models (Qwen2-VL with 7B and 72B parameters).
Figure 25: A case of comparison between separated and jointly modeling approaches with GPT-4o.

### v2

@@
-Multi-modal Retrieval Augmented Multi-modal Generation: A Benchmark, Evaluate Metrics and Strong Baselines
+Multi-modal Retrieval Augmented Multi-modal Generation: Datasets, Evaluation Metrics and Strong Baseliness
 Abstract
-This paper investigates an intriguing task of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG).
-This task requires foundation models to browse multi-modal web pages, with mixed text and images, and generate multi-modal responses for solving user queries, which exhibits better information density and readability.
-Given the early researching stage of M2RAG task, there is a lack of systematic studies and analysis.
-To fill this gap, we construct a benchmark for M2RAG task, equipped with a suite of text-modal metrics and multi-modal metrics to analyze the capabilities of existing foundation models.
-Besides, we also propose several effective methods for foundation models to accomplish this task, based on the comprehensive evaluation results on our benchmark.
-Extensive experimental results reveal several intriguing phenomena worth further research.
+We present a systematic investigation of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG), a novel task that enables foundation models to process multi-modal web content and generate multi-modal responses, which exhibits better information density and readability.
+Despite its potential impact, M2RAG remains understudied, lacking comprehensive analysis and high-quality data resources.
+To address this gap, we establish a comprehensive benchmark through a rigorous data curation pipeline, and employ text-modal metrics and multi-modal metrics based on foundation models for evaluation.
+We further propose several strategies for foundation models to process M2RAG effectively and construct a training set by filtering high-quality samples using designed metrics.
+Our extensive experiments demonstrate the reliability of our proposed metrics, a landscape of model performance within our designed strategies, and show that our fine-tuned 7B-8B models outperform the state-of-the-art GPT-4o model.
+Additionally, we perform fine-grained analyses across diverse domains and validate the effectiveness of our designs in data curation pipeline.
+All resources, including codes, datasets, and model weights, will be publicly released.
 […]
@@
 Multi-modal Retrieval Augmented Multi-modal Generation:
-A Benchmark, Evaluate Metrics and Strong Baselines
-Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
-Figure 1: A typical comparison between naive RAG and our proposed M2RAG. GPT-4o is used in this case.
-Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the dataset construction pipeline and Step 4 demonstrates the generation process.
+Datasets, Evaluation Metrics and Strong Baseliness
+Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3 Yu-Shi Zhu1, Tong Zhang1, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
+Figure 1: A typical comparison between naive RAG (upper) and our proposed M2RAG (lower). The generative model is GPT-4o in this case.
 1 Introduction
-Retrieval Augmented Generation (RAG) has been emerged as a promising technique to enhance the factual correctness of model-generated responses by grounding on facts in external knowledge databases Lan et al.
+Retrieval Augmented Generation (RAG) Lan et al.
 (2023); Sun et al.
 (2024); Li et al.
-(2022).
-Except for textual knowledge, multi-modal data, like images and videos, also contain diverse information, and it is natural to leverage these knowledge to answer user queries more accurately, leading to a concept named Multi-modal RAG or MuRAG Chen et al.
+(2022) and its multi-modal extensions  Chen et al.
 (2022); Joshi et al.
 (2024); Ye et al.
-(2024).
-Numerous works Riedler and Langer (2024) have demonstrated that RAG with multi-modal inputs outperforms single-modal RAG.
-This paper studies MuRAG, but we primarily focus on how to use foundation models to generate multi-modal responses for solving user queries, which we named M2RAG task, given multi-modal inputs with mixed layout of text and images.
-Numerous previous works Zhu et al.
+(2024) enhance foundation models by incorporating external knowledge and multi-modal data.
+While these methods improve response quality, they remain limited to textual outputs, which may fail to provide sufficient clarity in some scenarios, such as instructional guides or spatial reasoning.
+For instance, as shown in Figure 1, a text-only response explaining paper airplane folding steps may be difficult for users to follow without accompanying visual illustrations.
+As the saying goes, A single image is worth a thousand words.
+Visual elements can significantly enhance comprehension, especially in instructional and knowledge-intensive domains where textual descriptions alone may be insufficient.
+Motivated by this, the Multi-modal Retrieval-Augmented Multi-modal Generation (M2RAG) has been introduced  Zhu et al.
 (2024); Wu et al.
-(2023b); Li et al.
-(2024) have proven that multi-modal outputs effectively improve the information density and readability for users, embodying the principle—A single image is worth a thousand words.
-For example, Figure 1 compares plain text and multi-modal responses to the same user query about folding a paper airplane.
-While both responses provide similar textual explanations, the inclusion of instructional images in the multi-modal response significantly enhances readability.
-As a result, human readers can understand the instructions more quickly and efficiently, enabling them to successfully create a paper airplane with less effort.
-Despite the emerging interest in the M2RAG task, there is a lack of systematic analysis and evaluation, given its early stage of research development.
-To address this gap, we propose a benchmark designed to evaluate the capabilities of existing generative models, including Large Language Models (LLMs) and Multi-modal Large Language Models (MLLMs), in completing the M2RAG task.
-This benchmark evaluates the generated content from both text-modal and multi-modal perspectives, incorporating four metrics for each dimension, such as fluency, relevance, image coherence, and helpfulness.
-The benchmark dataset is constructed through a multi-step workflow, as illustrated in Steps 1-3 of Figure 2: (1) Query Collection: We gather a diverse set of user queries spanning ten distinct topics.
-For each query, the top-$k$ relevant multi-modal web pages are retrieved using the Google Search API; (2) Data Cleaning: A rigorous data cleaning pipeline is applied to these web pages to extract and filter high-quality text and images, which are then used to build a reliable multi-modal database for retrieval; (3) Retrieval: A suite of retrieval methods is developed to evaluate the relevance between multi-modal data elements and the user query.
-Each element is assigned a relevance score to guide its selection for further processing.
-Additionally, we propose two approaches for tackling this task, based on comprehensive studies using the constructed benchmark: (1) Single-stage Approach: This method generates a multi-modal response directly in a single turn, with images represented as placeholders within the response; (2) Multi-stage Approach: This method first generates a text response and then interleaves relevant images, finally refines the text response with retrieved images to ensure a more cohesive and enriched output.
-Extensive experimental results uncover numerous intriguing phenomena, providing valuable insights for future researches.
-For example, the multi-stage approach integrates more relevant images into responses, significantly outperforming the single-stage approach.
-Moreover, existing MLLMs demonstrate notable limitations in handling M2RAG tasks, performing 8% and 13% worse than LLMs under the single-stage and multi-stage approaches, respectively, which underscores substantial room for improvement.
-Importantly, compared to separate modeling approaches employed in prior works such as MuRAR Zhu et al.
-(2024), jointly modeling the relationships and interactions between images and text results in noticeable performance enhancements for the M2RAG task.
-These findings highlight the necessity of advancing joint modeling techniques to unlock the full potential of multi-modal generative models.
-Overall, our contributions are summarized as follows:
-  • Task Formulation and Benchmark: We introduce the M2RAG task and establish a comprehensive benchmark, alongside a suite of fine-grained text-modal and multi-modal metrics to evaluate the capabilities of existing models.
-  • Proposed Baselines: We propose several strong baseline methods for solving the M2RAG task, including single-stage and multi-stage approaches.
-  • Experimental Insights: Our experiments demonstrate that current LLMs outperform MLLMs significantly in M2RAG tasks.
-Additionally, decomposing this complex task into simpler sub-tasks further enhances performance.
-We also observe that relying solely on images in this task can result in poor performance for certain topics, suggesting that incorporating additional modalities could yield further improvements.
+(2023); Li et al.
+(2024), a novel task that retrieves and integrates multi-modal content to produce responses with a mixed layout of text and image.
+By integrating relevant images into textual responses, M2RAG enhances comprehension and usability across various application scenarios.
+As illustrated in Figure 1, M2RAG not only generates textual explanations for step-by-step folding a paper airplane but also seamlessly incorporates illustrative images where necessary, significantly improving user understanding.
+Compared to existing RAG methods, M2RAG introduces new challenges in multi-modal reasoning and synthesis.
+It requires foundation models to achieve a deeper understanding of multi-modal data, accurately capture content and cross-modal relationships, and generate coherent responses that seamlessly integrate text and images.
+However, given the early research stage, there is a lack of systematic analysis and high-quality data resources for the M2RAG task.
+To fill this gap, we make following contributions: (1) Benchmark: A comprehensive benchmark with 10 topics constructed by our proposed rigorous data curation pipeline; (2) Evaluation Metric: A suite of reliable text-only and multi-modal evaluation metrics based on Large Language Models (LLMs) and Multi-modal LLMs (MLLMs); (3) Generation Strategy: Two generation strategies for effectively tackling M2RAG task: single-stage and multi-stage approaches; and (4) Training Dataset: A training dataset constructed by filtering high-quality samples using our designed multi-modal evaluation metrics, which is used to improve the performance of 7B-8B LLMs and MLLMs.
+Based on our extensive experiments and automatic evaluation of advanced LLMs and MLLMs, we present four key findings: (1) Metric Reliability: The used evaluation metrics exhibit strong correlations with human judgments, even outperform inner-correlation of human annotators in evaluating image helpfulness dimension.
+This validates the reliability of our automatic evaluation framework; (2) Model Performance Analysis: Experimental results on 12 advanced LLMs and MLLMs provides valuable insights for M2RAG task.
+For example, LLMs consistently outperforms MLLMs, revealing MLLM’s limited multi-modal understanding and generation abilities.
+Besides, multi-stage strategy integrates more relevant images into responses, exhibiting better overall quality; (3) Effectiveness of Training Dataset: Fine-tuning 7B-8B LLMs and MLLMs on our curated dataset yields performance that exceeds GPT-4o, underscoring the quality and effectiveness of our training dataset; and (4) Data Curation Benefits: Ablation studies quantify cross-domain performance variations and verify the positive contributions of our designs in benchmark data curation pipeline, including detailed context for images and the inclusion of the auxiliary images.
+These observations and phenomena promote an in-depth understanding of M2RAG.
+We hope our data resources and these discoveries could spur future research in this field.
 2 Related Work
-2.1 Text-modal Retrieval Augmented Generation (RAG)
+Text-modal RAG
 RAG Li et al.
@@
 (2023).
-Recently, given the increasing capabilities of LLMs to effectively process any text inputs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al.
+Recently, given the increasing capabilities of LLMs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al.
 (2024).
-2.2 Multi-modal RAG
-Although typical RAG framework demonstrates great potential, it can only handle textual inputs and cannot utilize the rich information presented in multi-modal data such as images and videos Su et al.
+Multi-modal RAG
+Despite the great potential of RAG, it can only handle textual inputs and cannot utilize the rich information in multi-modal data such as images and videos Su et al.
 (2023).
-While earlier efforts to incorporate multi-modal inputs for text generation [^1] Most existing multi-modal works mainly focus on the visual modality.
-often relied on carefully designed frameworks Xiao et al.
+While earlier efforts to incorporate multi-modal inputs for text generation often relied on carefully designed frameworks Xiao et al.
 (2024), a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently Chen et al.
@@
 (2024).
-While these multi-modal models greatly expand the range of applications for LLMs, they remain limited to text-only outputs, restricting their ability to directly deliver rich, multi-modal information to users Zhu et al.
+While these multi-modal models greatly expand the range of LLMs’ applications, they remain limited to textual outputs, restricting their ability to deliver rich multi-modal responses to users Zhu et al.
 (2024).
-2.3 Multi-modal Generation
+Multi-modal Generation
 In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts Zhu et al.
@@
 Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT Wu et al.
-(2023b), TextBind Li et al.
+(2023), TextBind Li et al.
 (2024), Janus Wu et al.
@@
 To the best of our knowledge, MuRAR Zhu et al.
-(2024) is the most closely related work to our study.
-However, it falls short of modeling multi-modal content holistically.
-Specifically, it generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data.
-This approach differs significantly from the setup of our proposed M2RAG task, which also emphasizes the integration of multi-modal content in the input.
+(2024) is the most closely related works to our study.
+However, they falls short of modeling multi-modal content holistically.
+For example, MuRAR generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data.
+These approaches differs significantly from the setup of our proposed M2RAG task, which emphasizes the understanding the content and relationship among multi-modal input data.
 3 Task Formulation
-In M2RAG task, models answer each user query $Q$ with a multi-modal response based on retrieved multi-modal web pages.
-Suggest that multi-modal data resource $R=\{D_{1},\cdots,D_{n}\}$ represents $n$ retrieved multi-modal documents or web pages from the Internet.
-Each document $D_{i}$ consists of $m$ ordered elements $\{E_{i,1},E_{i,2},\cdots,E_{i,m}\}$, where each element $E_{i,j}$ can be either a textual or visual element.
-Elements are preserved in their original order within a document, ensuring that images are accompanied by their corresponding textual contexts.
-These contexts often provide rich descriptions or additional information that enhances the understanding and relevance of the images.
-With the above inputs, the M2RAG task needs several steps to get the multi-modal response.
-Specifically, the following procedures are essential: (1) In-Doc Retrieval, (2) Generation and (3) Multi-modal Element Incorporation.
+In M2RAG task, models generate a multi-modal response $r$ to each user query $Q$ by summarizing a retrieved multi-modal knowledge base $K=\{D_{1},\cdots,D_{n}\}$, consisting of $n$ multi-modal documents or web pages from the Internet.
+Each document $D_{i}$ consists of $m$ ordered elements $\{E_{i,1},E_{i,2},\cdots,E_{i,m}\}$, where each element $E_{i,j}$ can be either a text paragraph or an image.
+The original order of elements within a document is preserved to ensure images remain paired with their associated textual context, providing rich descriptions that enhance image comprehension.
+The process of generating the multi-modal response involves two key steps: In-Doc Retrieval and Generation.
 In-Doc Retrieval
-While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time increases significantly as input sequences grow longer Sun et al.
+While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time grows significantly with longer input sequences Sun et al.
 (2024).
-This challenge is more severe in the M2RAG task, where models must process both textual content and corresponding images simultaneously.
-To mitigate this issue, we first propose the In-Doc Retrieval process to collect relevant elements from $R$, which are used for next step multi-modal generation.
-As illustrated in Step 3 of Figure 2, we employ a retrieval model, denoted as $M_{R}$, to assess the relevance of each element $E_{i,j}$ in the dataset $R$ with respect to the user query $Q$, denoted as $M_{R}(Q,R)$.
-The top-$k$ most relevant elements are selected to form a refined multi-modal data resource, $R_{\text{In-Doc}}$.
-The refined resources serve as the inputs for generative models.
-Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input data for the generation process.
+This challenge is more severe in the M2RAG task, where models must handle extensive textual and visual data across multiple documents.
+To address this challenge, we introduce the In-Doc Retrieval, which selects the most relevant and useful elements from $K$ to reduce inference costs.
+As illustrated in Figure 2 (Step 3), a retrieval model $M_{R}$ assesses the relevance of each element $E_{i,j}$ in the knowledge base $K$ with respect to the user query $Q$: $M_{R}(Q,K)$.
+The top-$k$ most relevant elements are selected to form a refined and concise knowledge base, $K_{\text{In-Doc}}$.
+Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input information for the generation process.
+The implementation details of the retrieval model is provided in Section 4.1.3.
 Generation
-The refined multi-modal data resource $R_{\text{In-Doc}}$ is then processed by the generative model $M_{G}$ alongside the user query $Q$ to generate the final multi-modal response:
-$r=M_{G}(Q,R_{\text{In-Doc}})$ (1)
+The refined knowledge base $K_{\text{In-Doc}}$ is then processed by the generative model $M_{G}$ alongside the user query $Q$ to generate the final multi-modal response:
+$r=M_{G}(Q,K_{\text{In-Doc}})$ (1)
 Here, $r$ represents the final answer to the user query.
-Notably, if no visual elements are in $R_{\text{In-Doc}}$, the task defaults to the traditional RAG, since no multi-modal elements are utilized and generated.
-Multi-modal Element Incorporation
-Different from previous works that retrieve multi-modal data after text generation Zhu et al.
-(2024), we proposed a novel approach that can incorporate multi-modal data with text in a single-turn conversation.
-Specifically, images in the input and output are presented as placeholders.
-Each placeholder follows the following format: ![IMAGE_CONTENT](PSEUDO_URL).
-The pseudo URL identifies which image the model intends to include Wu et al.
-(2023a) in the output.
-For input images, the content is represented as tokenized image embeddings for MLLMs or as detailed image captions for LLMs.
-For output images, the image content is a concise caption based on the input image content which aids in the Chain-of-Thought (CoT) process Wei et al.
-(2022), assisting the model in understanding the image content or caption and placing it appropriately.
-4 Dataset Construction
-As shown in Step 1-3 of Figure 2, we elaborate three detailed steps to construct the benchmark dataset for evaluating the capabilities of foundation models in M2RAG task: (1) Query Collection, (2) Data Preparation and (3) Element Evaluation.
-4.1 Query Collection
+Notably, if no visual elements are in $K_{\text{In-Doc}}$, the task defaults to the traditional RAG, since no multi-modal elements are utilized and generated.
+Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline.
+Step 1-3 represent the benchmark data curation pipeline and Step 4 demonstrates our proposed generation strategies.
+4 Datasets and Methodology for M2RAG
+We describe details of our work: (1) Benchmark (Section 4.1); (2) Evaluation Metrics (Section 4.2); (3) Generation Strategy (Section 4.3); and (4) Training Dataset (Section 4.4).
+4.1 Benchmark Construction
+Our benchmark are established by running following steps: (1) Query Collection; (2) Data Preparation; and (3) Retrieval.
+4.1.1 Query Collection
 The queries are sourced from the ELI5 Fan et al.
 (2019) dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2.
-Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for providing comprehensive answers.
-After filtering, the remaining queries are classified to ensure a balanced dataset and lay the groundwork for further studies on the topic sensitivity of the proposed methods [^2] Please refer to Appendix A for more details..
-The classification process results in 11 distinct categories, as detailed in Table 1.
-20 queries are sampled in each category (other than “Others”), resulting in a final benchmark dataset containing 200 queries.
-ID Category Name Num. Samples
-1 Society & Culture 14182
-2 Science & Mathematics 14449
-3 Health 10931
-4 Education & Reference 776
-5 Computers & Internet 6999
-6 Sports 1974
-7 Business & Finance 3821
-8 Entertainment & Music 1946
-9 Family & Relationships 820
-10 Politics & Government 4638
-11 Others 451
-Table 1: Topics and volumes of the queries.
-4.2 Data Preparation
-All multi-modal data used for retrieval and generation are prepared during this step.
-Specifically, data for each query are processed through the following procedures: (1) Prepare Web Pages and Auxiliary Images, (2) Text Processing and (3) Image Processing.
-4.2.1 Prepare Web Pages and Auxiliary Images
-For each collected query, we first employ Google Search API [^3] https://developers.google.com/custom-search/v1 to get the URLs of web pages given the query, and then retrieve each web page containing a mix of text and images in markdown format using JINA AI Reader [^4] https://jina.ai/reader/.
-Our preliminary study reveals that the images within these web pages are sometimes inaccessible, low quality, or insufficient for effective multi-modal generation.
+Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for answers.
+The filtered queries are then classified into 11 categories to ensure dataset balance and support further investigation of the proposed methods’ topic sensitivity.
+[^1] See Appendix B for details.
+4.1.2 Data Preparation
+Subsequently, we crawl the multi-modal data related to each query and post-process them by conducting following steps (Step 2 of Figure 2): (1) Web Page and Auxiliary Image Crawling; (2) Text Processing; and (3) Image Processing.
+Web Page and Auxiliary Image Crawling
+For each collected query, we first employ Google Custom Search API [^2] https://developers.google.com/custom-search to get the URLs of web pages given the query, and then retrieve each web page containing a mix of text and images in markdown format using JINA AI Reader [^3] https://jina.ai/reader/.
+Our preliminary study reveals that the images within these web pages are sometimes inaccessible, of low quality, or insufficient for effective multi-modal generation.
 Therefore, we supplement the dataset with auxiliary images sourced from Google Image Search, using the user queries as search input to ensure relevance and enrich the visual content.
-Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset, as illustrated in Step 2 of Figure 2.
-4.2.2 Text Processing
-The raw web pages often contain a significant amount of irrelevant or redundant text, like the advertisements and web links, and it is essential to remove these data and produce concise and relevant content for generation.
-The text processing process is divided in to three steps: (1) Embedded Image Handling: image URLs are extracted from the web pages for image collection, and placeholders are inserted in their place within the text; (2) Text Cleaning: text snippets matching pre-defined patterns, such as web links, are removed to ensure cleaner and more relevant content; (3) Text Splitting: the cleaned text content is segmented into smaller pieces using the LangChain Chase (2022) text splitter, facilitating further selection and processing.
-4.2.3 Image Processing
-We conduct following steps to process images, ensuring they are of high quality and relevant to the user query: (1) Caching & Conversion: we download all the images with their URLs, and convert them into widely accepted formats such as JPG, PNG, GIF, or WEBP.
-Images that fail to be downloaded or converted are discarded; (2) Filtering: images that fall below a certain size threshold or receive a low similarity score with CLIP Radford et al.
-(2021) are removed.
-Such images are typically insufficiently informative, often consisting of icons, banners, or other non-representative visuals.
-(3) Deduplication: PHash algorithm-based deduplication Zauner (2010) is applied to remove duplicated or similar images, preventing redundant images for generation.
-4.3 Element Evaluation
-As described in Section 3, we need to perform In-Doc retrieval to select text and visual elements for efficient generation.
-To achieve this goal, the relevance between each element and the user query should be evaluated.
-Specifically, we utilized the DeepSeek-Chat DeepSeek-AI (2024) API for evaluating textual elements and MiniCPM-V-2.6 Yao et al.
-(2024) for assessing visual elements.
-Our preliminary manual annotations revealed that these methods demonstrate a high degree of consistency with human evaluations in measuring correlation, making them suitable for this task [^5] Please refer to Appendix B for more details..
-This step results in a score from 0 to 10 for each element, where higher scores indicate better relevancy.
-This can be used to replace the traditional embedding-based similarity score for retrieval.
-With the above procedures, a high-quality multi-modal RAG dataset with real-world user queries, retrieved web pages and images is constructed.
-Detailed statistical information of our constructed dataset can be found in Appendix C.
-5 Strong Baselines for M2RAG Task
-Finishing Step 1-3 of Figure 2, we introduce three strong baselines for Step 4 to accomplish our proposed M2RAG task.
-5.1 Separate Modeling
-Previous studies have approached text generation and image integration as distinct processes, exemplified by the MuRAR method Zhu et al.
+Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset.
+Text Processing
+Raw web pages often include irrelevant or redundant text, such as advertisements and web links, which must be removed to produce concise and relevant content for generation.
+The text processing pipeline consists of three steps: (1) Embedded Image Handling: Extract image URLs from web pages for image collection and replace them with placeholders in the text; (2) Text Cleaning: Remove text snippets matching predefined patterns, such as web links, to ensure cleaner and more relevant content; (3) Text Segmentation: Employ the text splitter Chase (2022) to split the cleaned text into smaller pieces, enabling efficient selection and processing.
+Image Processing
+We conduct following steps to process images to ensure they are of high quality and relevant to the user query: (1) Caching & Conversion: All images are downloaded using their URLs and converted into widely accepted formats such as JPG, PNG, GIF, or WEBP.
+Images that cannot be successfully downloaded or converted are discarded; (2) Filtering: images smaller than a certain threshold or with a low CLIP-based Radford et al.
+(2021) similarity score to the query text are removed.
+Such images often consist of non-representative visual contents, such as icons, banners, etc.
+(3) Deduplication: duplicate or highly similar images are removed using PHash Zauner (2010) algorithm.
+4.1.3 Element Evaluation
+As described in Section 3, we need to perform In-Doc retrieval to select top-$k$ relevant text and image elements for efficient generation.
+To achieve this goal, we utilize DeepSeek-Chat DeepSeek-AI (2024) and MiniCPM-V-2.6 Yao et al.
+(2024) for evaluation, as shown in Figure 2 (Step 3).
+Our preliminary manual annotations reveal that these methods demonstrate strong correlations with human evaluations in measuring correlation and significantly outperforms the traditional embedding-based models like CLIP Radford et al.
+(2021), thereby resulting in better quality of our benchmark [^4] Please refer to Appendix C for more details..
+In summary, given these data resources, we sample 75 queries for each category (other than Others), resulting in 750 queries for the benchmark dataset.
+4.2 Evaluation Metrics
+To examine the performance of LLMs and MLLMs on our proposed benchmark, we introduce four text-modal and four multi-modal metrics.
+Since the generated responses are open-ended, most evaluation metrics are implemented by prompting advanced GPT-4o model in a reference-free manner Lan et al.
+(2024b).
+Text-modal Metrics
+Following prior works Zhu et al.
+(2024), we evaluate the quality of textual generations using both linguistic and RAG metrics Es et al.
+(2023); Lan et al.
+(2024a): (1) Fluency assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al.
+(2022); Es et al.
+(2023); (2) Relevance evaluates how well the model-generated textual content aligns with the given user query Lan et al.
+(2022); Es et al.
+(2023); (3) Context Precision measures the proportion of relevant chunks in $K_{\text{In-Doc}}$ by analyzing the overlap of their key concepts Es et al.
 (2023).
-MuRAR operates in three sequential steps: (1) retrieving relevant plain text content and generating a text-based response; (2) identifying and incorporating appropriate visual elements into the text-based response; and (3) refining the text to seamlessly integrate the visual elements and their associated information.
-In our implementation, pre-trained embedding model Reimers and Gurevych (2019) is used to retrieve visual elements based on their surrounding text content or detailed image caption generated by GPT-4o.
-Since the separated method does not need to process multi-modal content, it is implemented by LLMs.
-5.2 Joint Modeling
-The separate approach overlooks the relationships between textual content and multi-modal data when generating text responses, leading to sub-optimal integration of multi-modal information.
-To address this limitation, we propose several baseline methods that jointly model text and images as a unified entity, capturing their rich relationships and interactive dynamics.
-Specifically, our approach can be categorized into single-stage and multi-stage methods, leveraging LLMs or MLLMs as backbone models.
-This design allows for a more holistic representation of multi-modal content, as depicted in Step 4 of Figure 2.
-Single- and Multi-stage Approaches
-Single-stage approach requires models to directly generated a multi-modal output with selected images in their corresponding placeholders, given all multi-modal content saved in a single prompt.
-Although the single-stage approach is straightforward, current open-source foundation models struggle in generating accurate image placeholders.
-Besides, it is a well-known limitation that most MLLMs may struggle with handling a large number of images simultaneously.
-Therefore, we propose a multi-stage approach to address the complexity of the single-stage approach, which involves three stages: (1) Text Response Generation: generate a plain text answer given the multi-modal input; (2) Image Interleaving: divide the text into several segments, and LLMs or MLLMs are prompted to determine which segment requires the inserted images for better readability; (3) Text Refinement: refine each segment by incorporating the information of selected images.
-Although multi-stage approach addresses the shortcomings of single-stage approach, their computational overhead for inference is much higher.
-LLM and MLLM Backbones
-Both of LLMs and MLLMs can be used to solve M2RAG task with single-stage or multi-stage approaches.
-Specifically, as introduced in Section 3, given the retrieved multi-modal data resource $R_{\text{In-Doc}}$, the prompt $P$ for LLMs and MLLMs could be constructed using a structured template $T$:
-$P=T(G,Q,R_{\text{In-Doc}})$ (2)
-where $G,Q$ denotes the textual task guidelines and user query, respectively.
-Then LLMs and MLLMs could be prompted with $P$ to solve M2RAG task: (1) For MLLMs, the encoded image tokens are placed at their original positions  Dubey et al.
-(2024); Wang et al.
-(2024a), reserving the relative position in the multi-modal web pages.
-Since the additional images do not belong to any web pages, their image tokens are placed in a distinct section in the prompt; (2) For LLMs, since they could not process images directly, the detailed captions of images in $R_{\text{In-Doc}}$ are provided, allowing LLMs to access the semantical content of images.
-6 Evaluation Metrics
-To evaluate the effectiveness of numerous models and methods, we design four text-modal metrics and four multi-modal metrics.
-Since the generated response are open-ended, these evaluation metrics are evaluated by prompting advanced GPT-4o model in a reference-free manner Lan et al.
-(2024).
-To the best of our knowledge, prompting GPT-4o for automatic evaluation is the most reliable automatic evaluation methods for open-ended RAG and multi-modal evaluation scenarios Es et al.
-(2023); Ku et al.
-(2024); Lu et al.
-(2023).
-The detailed prompts for our proposed text-only and multi-modal evaluation metrics can be found in Appendix D.
-The score ranges of all fine-grained text-only and multi-modal metrics are from 0 to 1, and scores close to 1 indicates better performance.
-6.1 Text-modal Metrics
-Following previous works Zhu et al.
-(2024), we utilize both linguistic and RAG metrics Es et al.
-(2023) to assess the quality of the generated content in the M2RAG task.
-These metrics ensure a comprehensive evaluation of the generated textual content.
-Fluency
-assesses the linguistic quality of model-generated text, ensuring the grammatical correctness, coherence, and readability Su et al. (2022); Es et al. (2023).
-Relevance
-assesses how pertinent the model-generated textual content is to the given user query Lan et al. (2022); Es et al. (2023).
-Context Precision
-evaluates the proportion of relevant chunks in the retrieved web pages by analyzing the overlap of their key concepts Es et al. (2023).
-Faithfulness
-measures the accuracy of the model-generated content in representing information from the original materials, focusing on factual alignment to avoid fabricated or misleading details Es et al.
+(4) Faithfulness measures the accuracy of generations in representing information from $K_{\text{In-Doc}}$, focusing on factual alignment to avoid fabricated or misleading details Es et al.
 (2023); Sun et al.
 (2024).
-6.2 Multi-modal Metrics
-In addition to text-only metrics, we also focus on the quality of model-generated multi-modal content from following five perspectives.
-Image Coherence
-examines the logical and coherent alignment of images with their surrounding text content, ensuring visual content enhances and complements the textual narrative Zhu et al. (2024).
-Image Helpfulness
-evaluates how effectively images contribute to the user’s comprehension of the text.
-This involves assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al.
-(2024).
-Image Reference
-verifies the appropriateness and alignment between images and their textual references. Inadequate or incorrect references of images lead to the confusion during reading.
-Image Recall
-measures whether the generated content includes highly relevant, informative, and important images.
-As described in Section 4.3, images are evaluated and scored during pre-processing, with selection based on these scores prior to generation.
-Consequently, the images input to generative models are generally relevant to the user query.
-However, issues such as semantic overlap or quality variations among images may arise, leading to the inclusion of sub-optimal images in the retrieved images.
-The goal of Image Recall is to ensure that no images other than the sub-optimal ones are excluded from the final answer.
-To achieve this, we prompt MLLMs to assess whether each image enhances human understanding of the content.
-To address potential semantic overlap, we include textual descriptions of images already present in the document and instruct the model to classify an image as negative if its content is significantly redundant compared to existing images.
-This process ensures that only diverse and meaningful images are retained.
-6.3 Overall Score Computation
-Except for these fine-grained evaluation dimensions, it is essential to summarize a final overall score reflecting the capabilities of foundation models in M2RAG task.
-Specifically, the average of four text-only metrics and five multi-modal metrics is computed as the overall score.
+Multi-modal Metrics
+Unlike evaluation metrics that focus solely on image quality Tu et al.
+(2024), our work evaluates the interplay between text and images, covering four key aspects: (1) Image Coherence examines the logical and coherent alignment of images with their surrounding text, ensuring that the visual content enhances and complements the textual narrative Zhu et al.
+(2024); (2) Image Helpfulness evaluates the contributions of images to the user’s understanding of the text, assessing whether visuals provide additional insights, clarify complex ideas, or support textual details Zhu et al.
+(2024); (3) Image Reference verifies the appropriateness and alignment between images and their textual references.
+Inadequate or incorrect references of images lead to the confusion during reading; (4) Image Recall measures the proportion of highly relevant, informative, and important images incorporated into the generations.
+These crucial images are evaluated and collected by Element Evaluation (Section 4.1.3).
+Overall Score Computation
+We compute the overall score by taking the average of all text-modal and multi-modal metrics, reflecting the overall capabilities of models.
+Scores are scaled to [0, 100], with higher values indicating superior performance.
+For additional details on the specific prompts used for each metric, please refer to Appendix E.
+4.3 Generation Strategy for M2RAG
+Previous work MuRAR Zhu et al.
+(2024) generates and integrates text and images independently.
+We refer to this approach as the separate strategy, which neglects interactions between multi-modal data during text generation.
+To address this limitation, we propose two joint modeling strategies that explicitly capture the relationships between text and images: the single-stage strategy and the multi-stage strategy, as illustrated in Step 4 of Figure 2.
+Single-stage
+requires the model to directly generate a multi-modal output by placing selected images into their corresponding placeholders, using all multi-modal content provided within a single prompt.
+Multi-stage
+addresses the common limitation of foundation models, which often struggle to process a large number of images simultaneously.
+It involves three stages: (1) Text Generation: generating a plain text answer based on the multi-modal input; (2) Image Interleaving: dividing the text into segments and prompting LLMs or MLLMs to identify which segments require image insertions for improved readability; (3) Text Refinement: refining each segment by incorporating information from the selected images.
+Both LLMs and MLLMs are able to handle M2RAG task with these two strategies.
+Specifically, LLMs and MLLMs get the prompt $P$ as input to handle M2RAG task, which is constructed using a structured template $T$:
+$P=T(G,Q,K_{\text{In-Doc}})$ (2)
+where $G,Q,K_{\text{In-Doc}}$ denotes the task guidelines, user query and knowledge base.
+In the prompt $P$, each image is represented using Markdown format: ![IMAGE_CONTENT](PSEUDO_URL): (1) For MLLMs, IMAGE_CONTENT refers to encoded embeddings of the image, with the entire placeholder positioned where the image originally appears in the context Dubey et al.
+(2024); Wang et al.
+(2024a), ensuring coherence between the image and the surrounding text; (2) For LLMs, we convert the image into a detailed textual description as IMAGE_CONTENT, enabling the LLMs to comprehend the semantic information of the image.
+PSEUDO_URL serves as the identifier for each input image.
+In the output of the single-stage approach, images are also represented using the aforementioned Markdown format, where PSEUDO_URL indicates the index of the selected image.
+4.4 Training Dataset Construction
+As another contribution to the research community, we construct a high-quality training dataset to enhance the performance of small-scale but efficient 7B-8B LLMs and MLLMs on the M2RAG task.
+Specifically, we first utilize the state-of-the-art GPT-4o with multi-stage strategy to construct 3K triplets of ($Q$, $K_{\text{In-Doc}}$, $r$).
+[^5] Experimental results in Section 5.2 suggest this setup leads to the best performance.
+Then, our proposed multi-modal metrics are used to filter the high-quality samples.
+We do not utilize the text-modal metrics, since the time cost of Context Precision and Faithfulness metrics on 3K long generations are huge [^6] Please refer to Section Limitations for more explanation., and we plan to improve the efficiency of text-modal evaluation metrics and update the training data in future work.
+Finally, the dataset consists of 1.6K instances.
+More implementation details and statistical information of our training dataset could be found in Appendix D.
 5 Experimental Results
-Model Type Approach Model Text-modal Metrics Multi-modal Metrics Overall
+We conduct the comprehensive evaluation results to demonstrate: (1) Reliability of our designed multi-modal automatic evaluation metrics (Section 5.1); (2) A landscape of advanced LLMs and MLLMs performance (Section 5.2); and (3) Effectiveness of our training dataset (Section 5.3).
+5.1 Validate Reliability of Evaluation Metrics
+Metric Human v.s. Max Min Avg.
+Image Coherence MLLM 0.56 0.36 0.48
+Human 0.55 0.46 0.51
+Image Helpfulness MLLM 0.67 0.48 0.57
+Human 0.61 0.47 0.53
+Image Reference MLLM 0.61 0.43 0.53
+Human 0.67 0.54 0.59
+Table 1: The Spearman’s correlation between MLLMs and human annotators. Max, Min and Avg. represents the maximum, minimum and average correlations with all human annotators except for themselves.
+Since the reliability of text-modal metrics have been proven in RAG scenarios Es et al.
+(2023), we mainly focus on validating the reliability of our proposed MLLM-based multi-modal metrics: (1) Image Coherence; (2) Image Helpfulness; and (3) Image Reference.
+In practice, we randomly select 200 samples from the benchmark labeled by each metric, and three annotators independently scored these samples using the same scoring criteria as the model.
+Specifically, we calculate the Spearman correlation coefficients between two scorers, each can be human annotators or MLLMs.
+Table 1 reveals that MLLMs achieve comparable performance to inner-correlation of human annotators.
+GPT-4o is even better than human average on helpfulness metric (0.57 > 0.53).
+These results indicate that our proposed multi-modal evaluation metrics are good proxy for human annotators.
+5.2 Overall Experimental Results
+Model Type Generation Strategy Model Text-modal Metrics Multi-modal Metrics Overall
 Flu. Rel. CP. Faith. Coher. Help. Ref. Recall
-LLMs Separate GPT-4o 0.80 0.82 0.69 0.89 0.53 0.44 0.43 0.43 0.63
-Single GPT-4o 0.79 0.81 0.76 0.91 0.69 0.57 0.31 0.57 0.68
-Llama-3.1-70B-Instruct 0.75 0.79 0.75 0.85 0.57 0.47 0.18 0.22 0.57
-Qwen2.5-72B-Instruct 0.79 0.83 0.77 0.87 0.62 0.52 0.35 0.52 0.66
-Llama-3.1-8B-Instruct 0.74 0.77 0.72 0.83 0.49 0.38 0.28 0.62 0.60
-Qwen2.5-7B-Instruct 0.71 0.73 0.73 0.91 0.52 0.41 0.22 0.63 0.61
-Average 0.75 0.79 0.75 0.87 0.56 0.45 0.27 0.51 0.62
-Multi GPT-4o 0.78 0.76 0.68 0.87 0.68 0.60 0.81 0.97 0.77
-Llama-3.1-70B-Instruct 0.73 0.75 0.72 0.84 0.68 0.55 0.77 0.97 0.75
-Qwen2.5-72B-Instruct 0.77 0.77 0.72 0.85 0.69 0.58 0.78 0.97 0.77
-Llama-3.1-8B-Instruct 0.72 0.73 0.72 0.82 0.66 0.55 0.75 0.93 0.74
-Qwen2.5-7B-Instruct 0.74 0.74 0.74 0.83 0.64 0.55 0.76 0.95 0.74
-Average 0.75 0.75 0.72 0.84 0.67 0.57 0.77 0.96 0.75
-MLLMs Single GPT-4o 0.78 0.82 0.76 0.90 0.64 0.53 0.24 0.60 0.66
-Llama-3.2-90B-V-Instruct 0.77 0.68 0.64 0.76 0.51 0.34 0.10 0.01 0.48
-Qwen2-VL-72B-Instruct 0.77 0.78 0.72 0.88 0.43 0.32 0.15 0.16 0.53
-Llama-3.2-11B-V-Instruct 0.77 0.64 0.60 0.68 0.26 0.21 0.02 0.02 0.40
-Qwen2-VL-7B-Instruct 0.66 0.64 0.71 0.83 0.39 0.30 0.23 0.15 0.49
-Average 0.75 0.71 0.69 0.81 0.53 0.43 0.22 0.19 0.54
-Multi GPT-4o 0.77 0.78 0.72 0.86 0.65 0.56 0.81 0.97 0.76
-Llama-3.2-90B-V-Instruct 0.72 0.70 0.67 0.74 0.55 0.43 0.62 0.94 0.67
-Qwen2-VL-72B-Instruct 0.75 0.74 0.71 0.80 0.63 0.52 0.73 0.94 0.73
-Llama-3.2-11B-V-Instruct 0.72 0.69 0.70 0.69 0.47 0.34 0.34 0.81 0.59
-Qwen2-VL-7B-Instruct 0.70 0.75 0.73 0.77 0.54 0.42 0.51 0.94 0.67
-Average 0.73 0.73 0.70 0.77 0.57 0.46 0.62 0.92 0.69
+Reasoner Single OpenAI o3-mini 80.8 89.5 91.9 90.0 73.3 63.2 64.6 79.1 79.1
+Multi OpenAI o3-mini 79.6 80.2 94.4 83.8 69.7 60.6 78.0 100.0 80.8
+LLMs Separate GPT-4o Zhu et al. (2024) 84.5 88.6 95.0 89.6 58.3 50.4 43.8 82.8 74.1
+Single GPT-4o 82.9 88.3 95.7 88.8 72.2 63.8 32.2 75.8 74.9
+DeepSeek-V3 81.2 87.9 91.9 87.4 63.6 52.7 36.8 87.3 73.6
+Llama-3.1-70B-Instruct 81.2 86.5 93.6 83.8 62.7 55.9 23.8 66.0 69.2
+Qwen2.5-72B-Instruct 82.5 89.2 95.8 87.2 66.3 59.6 36.6 78.6 74.5
+Llama-3.1-8B-Instruct 79.8 85.2 91.0 81.1 54.1 46.6 26.3 79.5 68.0
+Qwen2.5-7B-Instruct 74.2 79.4 88.9 82.1 55.9 48.4 22.2 80.0 66.4
+Average 80.3 86.1 92.8 85.1 62.5 54.5 29.6 77.8 71.1
+Multi GPT-4o 82.5 87.5 96.6 81.8 73.3 65.8 81.5 100.0 83.6
+DeepSeek-V3 80.5 83.0 95.0 85.9 70.0 63.1 79.1 100.0 82.1
+Llama-3.1-70B-Instruct 80.7 87.6 97.3 83.9 73.0 63.6 76.8 99.2 82.8
+Qwen2.5-72B-Instruct 83.8 88.6 98.4 83.1 73.4 64.5 78.5 99.9 83.8
+Llama-3.1-8B-Instruct 81.3 86.5 95.8 82.1 72.1 62.0 75.1 97.9 81.6
+Qwen2.5-7B-Instruct 82.8 87.0 95.3 78.5 71.3 62.4 77.9 99.0 81.8
+Average 81.9 86.7 96.4 82.5 72.2 63.6 78.1 99.3 82.6
+MLLMs Single GPT-4o 81.4 88.6 95.2 88.0 69.8 61.0 25.1 81.5 73.8
+Step-1o 84.1 87.4 94.9 90.7 65.9 53.8 44.6 84.7 75.8
+Llama-3.2-90B-V-Instruct 82.5 73.9 77.4 79.0 55.7 51.4 18.6 58.0 62.1
+Qwen2-VL-72B-Instruct 82.5 85.3 88.3 87.9 50.9 42.1 18.8 62.6 64.8
+Llama-3.2-11B-V-Instruct 83.0 71.4 73.0 70.7 37.8 20.6 1.1 58.0 51.9
+Qwen2-VL-7B-Instruct 69.0 70.6 79.0 72.9 44.7 39.3 23.6 61.5 57.6
+Average 80.4 79.5 84.6 81.5 54.1 44.7 22.0 67.7 64.3
+Multi GPT-4o 81.6 88.2 96.6 81.0 72.2 63.5 80.6 100.0 83.0
+Step-1o 80.5 82.5 96.7 89.8 71.6 63.7 81.6 97.6 83.0
+Llama-3.2-90B-V-Instruct 80.9 82.8 92.0 75.5 60.5 51.0 62.7 98.1 75.4
+Qwen2-VL-72B-Instruct 83.0 86.5 94.3 79.0 68.0 58.8 71.2 98.8 79.9
+Llama-3.2-11B-V-Instruct 82.8 80.7 87.8 69.9 52.7 42.7 34.4 86.2 67.1
+Qwen2-VL-7B-Instruct 80.6 86.7 93.1 79.2 60.3 49.1 49.9 96.0 74.4
+Average 81.6 84.5 93.4 79.1 64.2 54.8 63.4 96.1 77.1
+Fine-tuned Models Single Llama-3.1-8B-Instruct 74.9 76.5 96.1 82.4 66.3 60.6 75.2 97.0 78.6
+Qwen2.5-7B-Instruct 75.3 77.6 94.9 80.6 66.0 58.8 73.9 87.6 76.9
+Qwen2-VL-7B-Instruct 57.2 74.3 91.4 71.4 65.4 57.4 73.5 95.0 73.2
 Table 2: The overall experiment results on M2RAG task.
@@
 Coher., Help., Ref., and Recall represent image coherence, helpfulness, reference and recall, respectively.
-The highest scores for each group are highlighted in bold, and the highest average scores for 4 groups are highlighted in red.
-Using the proposed baseline methods and evaluation metrics, we conduct comprehensive experiments on the benchmark dataset with a range of representative models, including GPT-4o, the Llama-3.1 series, Qwen2.5 series, Llama-3.2-Vision series, and Qwen2-VL series [^6] Please refer to Appendix E for more details..
-In this section, we analyze the experimental results presented in Table 2 from the following perspectives.
+The highest scores for each group are highlighted in bold, and the highest scores for all settings are highlighted in red.
+We conduct comprehensive experiments on our benchmark with 12 representative models, including OpenAI o3-mini, GPT-4o, DeepSeek-V3, Step-1o, Llama-3.1 & Llama-3.2-Vision series, Qwen2.5 & Qwen2-VL series (see Appendix F for details).
+Due to the high evaluation cost, we sample 200 queries from the benchmark dataset by category to compare the performance of all models.
+The results are shown in Table 2 , and we can observe several findings.
 The Most Capable Models
-As shown in Table 2, it can be found that the most capable open-source and closed-source foundation models for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o (0.77), respectively, under the experimental setup of multi-stage approach with LLM backbones.
-Although the performance gap between these two models are negligible with single-stage approach, there is still a noticeable gap when it comes to single-stage approach(0.68 > 0.66).
-For MLLM backbones, GPT-4o significantly outperforms all closed-source models.
-For example, Qwen2-VL-72B-Instruct is much worse than GPT-4o by using the single-stage approach (0.53 < 0.66).
-Besides, there is still significant room for improvement in the performance of the most capable GPT-4o model (0.77 < 1.0) in the M2RAG task.
+Table 2 indicates that the most capable open-source and closed-source LLMs for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o, using a multi-stage strategy with LLM backbones.
+Among MLLM backbones, GPT-4o and Step-1o outperform all open-source models.
+For example, Step-1o surpassing Qwen2-VL-72B-Instruct by a large margin (83.0 > 79.9).
+The reasoner OpenAI o3-mini significantly outperforms other models in the single-stage approach.
+Despite these strong performances, there remains significant room for improvement, as even the best-performing model falls short of the ideal score (83.8 < 100.0).
 Scaling Phenomena
-Significant scaling phenomena are observed in most experimental setups, with model performance improving as model size increases.
-However, a notable exception is found in the single-stage approach with LLMs.
-For instance, Llama-3.1-8B-Instruct slightly outperforms Llama-3.1-70B-Instruct (0.60 > 0.57) under this approach.
-Rigorous human annotation reveals that this issue arises because larger models tend to select fewer images in the single-stage approach.
-It should be noted that many of the excluded images are beneficial for the response, and their omission results in performance degradation.
-For example, the Image Recall scores of Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct are significantly lower than those of Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct (0.22 < 0.62 and 0.52 < 0.63, respectively).
-This phenomenon highlights a limitation in the image selection behavior of larger models, which warrants further investigation in future work.
-Joint Modeling versus Separate Modeling
-It can be observed that the separate modeling method is much worse than all of our proposed approaches.
-Specifically, our worst approach, i.e., single-stage with MLLMs (GPT-4o) still outperforms it (0.66 > 0.63).
-This phenomenon demonstrates that jointly modeling the relationship and interactive information between multi-modal data is crucial for the M2RAG task.
-Comparison among Approaches
-As shown in the “Average” rows in Table 2, it can be found that the multi-stage approach significantly outperforms the single-stage approaches for LLMs models (0.75 > 0.62) and MLLMs models (0.69 > 0.54).
-Besides, multi-stage approach introduce more relevant images, enhancing the information density and readability of the generated answers.
-Specifically, the Image Recall and Helpfulness fine-grained scores of multi-stage approaches are higher than the single-stage approach.
-Moreover, the LLMs significantly outperforms the MLLMs in M2RAG tasks, no matter the single-stage or multi-stage is used.
-For example, the average overall score of LLMs are better than MLLMs (0.71 > 0.64).
-This observation demonstrates that the capabilities of existing MLLMs are not yet capable enough to accomplish the complex M2RAG tasks.
-Model Approach Prompt Completion Total
-LLMs Single 7797 767 8563
-Multi 16947 2252 19199
-MLLMs Single 11454 553 12007
-Multi 24078 1707 25785
-Table 3: Average token usage for different approaches.
-Furthermore, we also compute the average cost, i.e., the average prompt and completion tokens of GPT-4o, in Table 3, and we could summarize two conclusions: (1) the cost multi-stage approach is over 2 times that of single-stage; (2) the cost of MLLMs is approximately 1.3-1.4 times that of LLMs.
-Therefore, considering that LLMs generally performs better than MLLMs in M2RAG, using single-stage with LLMs as backbone model is currently the most cost-effective approach.
+A clear scaling trend is observed, where larger open-source models generally yield better results, except in the single-stage approach with LLMs, where Llama-3.1-8B-Instruct surprisingly outperforms Llama-3.1-70B-Instruct in Image Recall (79.5 > 66.0).
+This discrepancy arises because larger models tend to select fewer images, some of which are beneficial, leading to degraded performance.
+Comparison between Separate and Joint Modeling
+Regarding modeling approaches, separate modeling Zhu et al.
+(2024) proves significantly weaker than joint modeling, with even the worst-performing joint approach (GPT-4o, 74.9) surpassing separate modeling (74.1), highlighting the importance of modeling multi-modal interactions.
+Comparison among Strategies
+Multi-stage approaches consistently outperform single-stage methods for both LLMs (82.6 > 71.1) and MLLMs (77.1 > 64.3), as they introduce more relevant images, improving information density and readability.
+Additionally, open-source LLMs significantly outperform MLLMs across both modeling paradigms, with an overall score advantage (76.8 > 70.7), indicating that current open-source MLLMs still struggle with the complexity of M2RAG tasks.
+Notably, while being constrained by evaluation costs, to investigate whether these findings remain valid as the test data scale increases, we selected several representative models and evaluated them on all 750 queries.
+The results, presented in Appendix I, confirm that the findings remain valid.
+5.3 Improvements from Fine-tuning
+We further conduct the supervised fine-tuning to improve the 7B-8B LLMs and MLLMs capabilities in M2RAG task.
+Please refer to Appendix H for more fine-tuning implementation details.
+The evaluation results are illustrated in Table 2.
+It can be observed that our fine-tuned LLMs and MLLMs exhibit significant improvements on all multi-modal evaluation metrics and the overall score.
+For example, the average improvements on the overall score is 19.2%.
+Since the text-modal metrics are not utilized for constructing training dataset, our fine-tuned models underperforms the baselines on some textual evaluation metrics, particularly Fluency.
 6 Analysis
-In this section, we answer the following three important questions about our proposed approaches.
-8.1 Topic is a Key Factor Affecting Performance
+In this section, we present more analysis on our benchmark by answering several questions.
+6.1 Does Topic Affect Performance?
 The average overall scores of all evaluated models, as presented in Figure 3, reveal notable variations in performance across different topics.
-For instance, the most challenging topic is “Business & Finance”, which achieves the lowest overall score, whereas “Society & Culture” and “Sports” are significantly easier topics by comparison.
+For instance, the most challenging topic is Politics & Government, which achieves the lowest overall score, whereas Society & Culture and Sports are significantly easier topics by comparison.
 Figure 3: Average overall score across 10 topics.
@@
 Specifically, images are more prevalent and useful in less technical fields, such as sports and culture.
-In contrast, fields like business and finance often require data from additional modalities, such as tables and charts, which are less readily integrated.
-Moreover, theoretical fields like politics may not rely heavily on multi-modal information for comprehensive responses.
-To achieve balanced performance across topics, a topic-sensitive approach should be developed to adaptively incorporate appropriate modalities for response synthesis, ensuring that the specific requirements of each topic are effectively addressed.
-8.2 Caption Quality is Important for Image Integration
-As described in Section 5.2, images are converted into captions for use by LLMs.
-To ensure a comparable level of information between LLMs and MLLMs, the captions are crafted to be as detailed as possible.
-Additionally, textual contexts associated with the images are included during the captioning process to enhance consistency.
-In this section, we investigate the impact of removing these elements.
-Figure 4: Comparison of performance with original and ablation settings.
-As shown in Figure 4, performance declines noticeably with the removal of details and contextual information, particularly in the metrics of Image Coherence and Image Reference.
-The absence of contextual information in captions leads to a significant drop in Image Recall, highlighting the critical role of context in guiding the image selection process of LLMs.
-Beyond these key observations, the ablation settings show minor fluctuations in other metrics.
-Overall, incorporating context during captioning is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
-8.3 Contributions of Auxiliary Images
+In contrast, fields like politics and finance often require more accurate data from additional modalities, such as tables and charts, which are less readily integrated.
+These phenomena suggest that future work needs to carefully implement strategies to control the proportion of included image across different topics.
+6.2 Is Description Quality Important?
+As described in Section 4.3, to provide the LLMs with rich and accurate image content, each image is transcribed into a detailed textual description, with consideration of its original contextual information.
+In this subsection, we conduct ablation studies to investigate the contributions of these information.
+Table 3 reveals that model’s performance significantly declines by removing contextual information, particularly in the text-modal and multi-modal precision metrics.
+The absence of details in descriptions leads to a notable drop in Image Recall, highlighting the critical role of details in guiding image selection of LLMs.
+Overall, incorporating context is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
+Ablation
+Text
+Avg.
+Multi-Modal
+Precision Avg.
+Recall
+Score
+Overall
+Score
+Original 88.9 56.1 75.8 74.9
+w/o Context 88.5 51.4 76.8 73.1
+w/o Detail 88.8 54.4 73.5 74.0
+w/o Aux. Image 88.7 52.4 - -
+Table 3: Comparison of GPT-4o experiment and ablation study results.
+Text Avg.
+and Multi-Modal Precision Avg.
+represent the average values of four text-modal metrics and three multi-modal precision metrics (except for Recall), respectively.
+Recall and Overall scores for w/o Aux.
+Image is not reported due to changes in the images within the knowledge base.
+6.3 Are Auxiliary Images Helpful?
 Auxiliary images are collected for each query to address cases where highly relevant images are absent from the retrieved web pages.
-Removing these images leads to a significant decline in generative performance (Figure 4).
-For instance, Image Recall drops sharply without auxiliary images, suggesting that LLMs preferentially include auxiliary images over those sourced from web pages when constructing the generated document.
-Additionally, the decline in Image Coherence and Image Helpfulness metrics highlights the superior quality of auxiliary images compared to web-sourced images.
-Interestingly, the Image Reference metric shows a notable increase, likely due to stronger connections between each remaining image and its context.
+As illustrated in last row of Table 3 (w/o Aux.
+Image), removing these auxiliary images leads to a significant decline in generative performance in multi-modal metrics.
+This observation reflects the challenges in effectively integrating auxiliary images that lack contextual information with the textual content.
 Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
@@
 Extensive experimental results demonstrate several intriguing phenomena, facilitating the future research in this field.
-10 Limitations
+Limitations
 Although our proposed methods outperform some existing approaches for the M2RAG task, there are still several limitations.
-Topic Restrictions
-Certain topics, such as “Business & Finance” or “Politics & Government”, are not well-suited for visual representation.
-For these domains, tables, charts, or other formats may be more appropriate than images.
-The absence of these modalities limits the applicability of our approach to these areas.
-Content Availability
-The lack of extensive public discussion or data in certain specialized fields presents challenges for retrieving abundant multi-modal documents effectively.
-This limitation reduces the comprehensiveness of our analysis in underrepresented domains.
+More Modalities
+Our work mainly focus on the text and image modalities.
+In the real-world scenarios, information in other modalities are also crucial, such as speech and video.
+In the future, we plan to extend our proposed benchmark and training datasets to more modalities, even the mixed modalities.
 Evaluation Metrics
-The proposed evaluation metrics are limited and may fail to capture all aspects of M2RAG evaluation.
+The proposed evaluation metrics might fail to capture all potential aspects of M2RAG evaluation.
 For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer.
-However, using the LLM-as-a-Judge Zheng et al.
-(2023) approach, pairwise comparisons are resource-intensive, particularly for large-scale evaluations.
-Furthermore, their high cost and impracticality make them unsuitable for benchmarking, necessitating reliance on fixed scoring metrics that may not adequately address all edge cases.
-Appendix A Query Collection
+Besides, due to the requirement of using LLMs and MLLMs to evaluate the quality of long-text multi-modal content across 8 dimensions [^7] Please refer to Table 6 for more details about evaluation cost., this leads to huge evaluation costs in large-scale evaluation.
+To balance the affordability and reliability of experiments on our proposed benchmark, we introduce a quantity-first approach with 750 evaluation samples in this paper.
+In future work, we plan to distill and build efficient text-modal and multi-modal metrics to support larger-scale benchmark testing.
+Appendix A Ethical Considerations
+Most of the task inputs in our benchmark and training dataset are sourced from publicly available datasets, ensuring that they pose no harm to individuals or groups.
+Furthermore, the text generated by large language models (LLMs) and multi-modal language models (MLLMs) is carefully curated and processed by human annotators to safeguard privacy and confidentiality.
+No personally identifiable information (PII) is included.
+However, it is important to note that texts from the ELI5 dataset Fan et al.
+(2019) and multi-modal documents retrieved via Google Custom Search may contain harmful content or hate speech.
+Despite these potential risks, it is crucial to disclose the full scope of this research, as materials from ELI5 and Google Custom Search have been extensively used in safety research within the community.
+All annotators are compensated fairly, with an hourly wage of approximately \$4.76 USD, which exceeds the average hourly wage of \$3.13 USD on Amazon Mechanical Turk Hara et al.
+(2018).
+Appendix B Query Collection
 To simulate the real-world user query solving problem, we collect diverse and high-quality user queries from the ELI5 dataset Fan et al.
@@
 The prompt template for query classification is illustrated in Figure 6.
-Appendix B Retrieval of Collected Elements
-We replaced traditional embedding-based retrieval methods with LLM-based and MLLM-based approaches to enhance the retrieval process.
+ID Category Name Num. Samples
+1 Society & Culture 14182
+2 Science & Mathematics 14449
+3 Health 10931
+4 Education & Reference 776
+5 Computers & Internet 6999
+6 Sports 1974
+7 Business & Finance 3821
+8 Entertainment & Music 1946
+9 Family & Relationships 820
+10 Politics & Government 4638
+11 Others 451
+Table 4: Topics and volumes of the queries.
+Appendix C Retrieval of Collected Elements
+We replaced traditional embedding-based retrieval methods with LLM-based and MLLM-based approaches to improve the retrieval performance.
 Embedding-based methods typically prioritize semantic similarity but often overlook whether candidate elements provide substantial information relevant to the target question.
@@
 This evaluation is conducted by prompting LLMs and MLLMs for textual and visual content, respectively.
-The prompt templates used for these evaluations are illustrated in Figures 8 and 9.
+The prompt templates used for these evaluations are illustrated in Figures 7 and 8.
+This step results in a score from 0 to 10 for each element, where higher scores indicate better relevancy.
 # Task Description
@@
 Figure 8: Prompt template for the retrieval of visual elements.
-Appendix C Dataset Statistics
-C.1 Basic Statistics
+Appendix D Dataset Statistics
+D.1 Basic Statistics
 The statistical information about our benchmark are shown in Table 5.
-Item Value
-Queries 200
-Web pages per query 6.4
-Text elements per web page 7.2
-Images per query 9.3
-Auxiliary image per query 2.6
-Average score of text elements 1.1
-Average score of image elements 4.2
-Average auxiliary image score 3.6
+Item Range Value
+Query benchmark 750
+training 1605
+Web Page per query 9.8
+Image 12.0
+Aux. Image 3.0
+Text Element per web page 25.7
+Image 1.2
+Avg. Element Score text 1.5
+image 3.1
+aux. image 3.8
 Table 5: Statistical information of our benchmark.
-C.2 Data Distribution
+D.2 Data Distribution
 We list the distributions of some key items:
@@
 Figure 15: Score of auxiliary images
-Appendix D Details of Evaluation Metrics
+Appendix E Details of Evaluation Metrics
+E.1 Prompt Templates
 Apart from Context Precision and Faithfulness which are implemented with RAGAS, other metrics are implemented by prompting GPT-4o with the following prompt templates:
@@
   • Image Reference: as illustrated in Figure 20;
-  • Image Recall: as illustrated in Figure 22.
 # Task Description
@@
 # Task Description
-You are a multi-modal document evaluation assistant.
+You are a multimodal document evaluation assistant.
 You will receive an image and its textual context written in markdown.
@@
 markdown
-{context_below} ‘‘‘ # Scoring Criteria
-Assess how coherent the image is in relation to its accompanying text, assigning a score from 0 to 10.
-A higher score indicates stronger coherence between the image and the text.
-Be precise when assigning the score.
-- A score from 0-3 means that the image is minimally or not at all coherent with the text.
-- A score from 4-6 indicates that the image shows some coherence with the text but may include unrelated elements.
-- A score from 7-9 indicates that the image is highly coherent with the text.
-- A score of 10 indicates perfect coherence, where the image completely corresponds with and enhances the text.
-Be rigorous and discerning when assigning your score.
+{context_below} ‘‘‘ # Scoring Criteria and Examples
+Evaluate how well the image complements its accompanying text and assign a score from 0 to 10. A higher score reflects stronger coherence between the image and the text. Be precise in your evaluation.
+- Score 0-3: Minimal or no coherence between the image and text.
+- Example 1: The image shows a video game controller, but the text discusses supply and demand in virtual goods on platforms like Steam.
+These topics are unrelated, and the image adds no context to the discussion of virtual goods.
+The score should be 0-3.
+- Example 2: The image depicts a luxury car, but the text is about the environmental impact of electric vehicles.
+While both relate to automobiles, the image doesn’t contribute to the discussion of electric vehicles’ environmental effects.
+Thus, it merits a score of 0-3.
+- Example 3: A photo of a pizza is shown, but the text is a deep dive into the history of ancient Greece.
+There is no connection between the image and the historical content, making the coherence score 0-3.
+- Score 4-6: Some coherence, but with unrelated elements.
+- Example 1: The image shows various components like a remote control, cables, and a media player, which are related to the media player mentioned in the text.
+However, the context focuses on video formats, which is not addressed by the image.
+The coherence score should be between 4 and 6.
+- Example 2: The image shows graduates from Harvard University, and the text is about the cost of university education and the perceived benefits of attending prestigious institutions.
+While the image relates to the general topic of education, it doesn’t directly support the discussion about the financial costs or value of education, meriting a score of 4-6.
+- Example 3: The image of a crowded subway is referenced in the text discussing urbanization and its effects on public transport.
+However, the text focuses more on the environmental impact of urbanization rather than directly linking to the image of subway crowds, leading to a coherence score of 4-6.
+- Score 7-9: High coherence, with the image closely aligning with the text.
+- Example 1: The image is a photo of a video game controller, and the text describes features of the latest gaming console.
+The image directly aligns with the content, visually representing the technology described in the text, earning a score of 7-9.
+- Example 2: The image depicts a close-up of a laptop keyboard, and the text discusses advancements in laptop design.
+The image complements the text well, helping the reader understand the specific design features mentioned, justifying a score of 7-9.
+- Example 3: A chart showing sales data for various smartphone models is paired with text explaining smartphone market trends.
+The image is directly relevant to the discussion of market share and trends, aiding comprehension, and warrants a score of 7-9.
+- Score 10: Perfect coherence, where the image completely corresponds to and enhances the text.
+- Example 1: The image is a detailed picture of a video game controller, and the text provides a comprehensive review of the controller’s features, design, and performance.
+The image fully complements the text, making the reader’s understanding more complete and vivid, deserving a score of 10.
+- Example 2: The image shows a close-up of a public toilet seat, which is relevant to the context below that discusses the design differences between public and home toilet seats.
+The image is completely aligned with the text, directly aiding the discussion.
+The score should be 10.
+- Example 3: The image is an anatomical diagram of the human heart, and the text explains how blood circulates through the body.
+The visual directly supports the explanation, making the text easier to understand and reinforcing key points, justifying a score of 10.
+Note: The image needs to be coherent with the context above OR below it, not necessarily both.
+Be rigorous and objective in your assessment.
 # Output Instructions
@@
 # Task Description
-You are a multi-modal document evaluation assistant.
+You are a multimodal document evaluation assistant.
 You will receive an image and its textual context written in markdown.
@@
 markdown
-{context_below} ‘‘‘ # Scoring Criteria
-Evaluate how well the image helps human readers understand the content of its accompanying text, assigning a score from 0 to 10.
-A higher score indicates that the image significantly enhances comprehension of the text.
-Be as precise as possible when assigning the score.
-- 0-3: The image is minimally or not at all helpful for comprehension.
-- 4-6: The image provides some helpful context or information but may contain extraneous or less relevant details.
-- 7-9: The image is highly helpful in enabling comprehension of the text.
-- 10: The image is perfectly helpful and greatly enhances and clarifies the information provided in the text.
-Be rigorous and discerning when assigning your score.
+{context_below} ‘‘‘ # Scoring Criteria and Examples
+Assess the image’s helpfulness in improving comprehension of the text, assigning a score from 0 to 10.
+A higher score indicates the image significantly enhances understanding.
+Be precise and rigorous when assigning the score.
+- Score 0-3: The image is minimally or not at all helpful for comprehension.
+- Example 1: The image depicts a book cover titled "When God Asks You to Do Something You Don’t Want to Do," with a background of crashing ocean waves.
+The text is about jury members addressing trial observations.
+The image is irrelevant and unhelpful, meriting a score of 0-3.
+- Example 2: The image shows a close-up of banknotes (euro), arranged in a fan-like pattern.
+While the image is related to the concept of money, it does not directly enhance the comprehension of the text, which discusses the exchange rate of different currencies.
+Thus, it is not helpful for comprehension, earning a score of 0-3.
+- Example 3: The image is of a person looking at a blank computer screen, with no connection to the text discussing how the internet has influenced modern literature.
+The image does not provide any useful context or additional understanding of the text, so it warrants a score of 0-3.
+- Score 4-6: The image provides some helpful context but may include extraneous or less relevant details.
+- Example 1: The image shows a man at a desk, shocked at his laptop, with colorful text and symbols in the background.
+The text discusses how hackers are caught.
+While the image is thematically related, it is not directly helpful in explaining the topic of cyber security, earning a score of 4-6.
+- Example 2: The image shows a car with a vented hood, which relates to the text discussing the aerodynamic and cooling benefits of such designs in mid-engine cars.
+However, the image doesn’t address specific details, like material considerations or the pressure zones discussed in the text, meaning it provides partial context but not a complete understanding.
+Therefore, it should be assigned a score of 4-6.
+- Example 3: The image of a group of college students at a campus event is referenced in the text discussing the benefits of a liberal arts education.
+While it offers a visual representation of student life, it doesn’t fully enhance the understanding of the specific academic benefits described, so it earns a score of 4-6.
+- Score 7-9: The image is highly helpful in enhancing comprehension of the text.
+- Example 1: The image features a table comparing popular browsers and their operating systems, such as Edge, Safari, and Chrome.
+The text discusses why Internet Explorer (IE) is disliked.
+The image is highly relevant, providing a clear visual context that directly supports the reader’s understanding of the browser landscape, justifying a score of 7-9.
+- Example 2: The image is a close-up of a laptop keyboard, with highlighted keys.
+The text discusses the evolution of laptop designs, and the image directly supports the explanation, helping the reader understand the physical changes over time, which justifies a score of 7-9.
+- Example 3: The image shows the front page of a scientific journal, with headlines and an image illustrating the main topic.
+The text explains the content of the journal in detail, and the image directly helps readers better visualize the article’s focus, justifying a score of 7-9.
+- Score 10: The image perfectly enhances and clarifies the text.
+- Example 1: The image is a detailed diagram of the human digestive system, showing organs like the stomach and intestines.
+The text explains the process of digestion, and the diagram complements this explanation perfectly, enhancing the reader’s understanding of the content.
+The image and text work together seamlessly, earning a score of 10.
+- Example 2: The image shows a close-up of a public toilet seat, which is relevant to the context below that discusses the design differences between public and home toilet seats.
+The image directly complements the text, providing the exact visual context needed to understand the design differences.
+The score should be 10.
+- Example 3: The image is a map highlighting the trade routes discussed in the text about ancient civilizations.
+The map directly enhances the comprehension of the trade dynamics and locations mentioned in the text, offering a visual aid that clarifies the complex subject matter, thus earning a score of 10.
+Note: The image needs to be helpful for the context above OR below it, not necessarily both.
+Be rigorous and objective in your assessment.
 # Output Instructions
@@
 # Task Description
-You are a multi-modal document quality assessment assistant.
+You are a multimodal document quality assessment assistant.
 You will receive an image and its accompanying textual context, formatted in markdown.
@@
 {context_below} ‘‘‘ # Scoring Criteria
-Evaluate the extent to which the image is referenced or explained in the text:
-- 0: The image is not mentioned or referenced in the context.
-- 1-3: The image is referenced implicitly, and the reference is improper or incorrect.
-- 4-6: The image is referenced explicitly but in an improper manner, or it is referenced implicitly.
-- 7-9: The image is referenced explicitly, with the reference being generally proper and correct.
-- 10: The image is referenced explicitly, with the placement and explanation being completely proper and correct.
-Be rigorous and discerning when assigning your score.
+Determine how well the image is referenced or explained in the surrounding text, assigning a score from 0 to 10:
+- Score 0: The image is not mentioned or referenced in the text.
+- Example 1: The image is present, but there is no mention or acknowledgment of it in either the context above or below.
+- Example 2: The image shows a scenic mountain view, but the text is entirely about office workflows, with no reference to or explanation of the image.
+- Example 3: The image is of a fruit basket, but the text is about urban planning, without any reference to the image or its context.
+- Score 1-3: The image is referenced implicitly, but the reference is inapparent, improper, or incorrect.
+- Example 1: The text discusses the Millennium Falcon’s hyperdrive system and kyber crystal energy field, which are depicted in the image.
+However, the image is not explicitly mentioned, and the connection is weak.
+The score is 1-3.
+- Example 2: The text mentions "the latest in gaming technology," and the image shows a futuristic gaming console.
+The image is not directly referenced, and the connection to the text is not clearly established.
+The score is 1-3.
+- Example 3: The text talks about space exploration and mentions rockets, but the image is a close-up of a satellite.
+The connection between the satellite and the discussion of space exploration is implied but not directly referenced.
+Thus, the score is 1-3.
+- Score 4-6: The image is referenced implicitly or explicitly, but the reference is improper or partially relevant.
+- Example 1: The image shows a skier, and the text explicitly references it, but the discussion focuses on ski races, which is only loosely related to the skier’s role.
+The reference is valid but not fully accurate, so the score is 4-6.
+- Example 2: The image shows a tree with roots exposed, and the text references “the importance of deep roots” in discussing personal growth.
+While the image supports the metaphor, the connection is not fully fleshed out, leading to a score of 4-6.
+- Example 3: The image is of a sunset, and the text discusses the environmental impact of light pollution.
+While the sunset relates to the topic of light, it does not directly contribute to the explanation of light pollution, resulting in a score of 4-6.
+- Score 7-9: The image is explicitly referenced in a generally proper and correct manner.
+- Example 1: The image depicts a box of chips and is explicitly mentioned in the text at an appropriate point.
+While the reference is accurate, it might feel a little stiff, leading to a score of 7-9.
+- Example 2: The text discusses the structural components of a building, and the image shows a detailed diagram of the beams and supports.
+The reference is clear and mostly accurate, with a slight stiffness in phrasing, justifying a score of 7-9.
+- Example 3: The image depicts a city map, and the text discusses the layout of different districts.
+The map is explicitly referenced and properly aids the understanding of the text, but the reference could be smoother.
+The score should be 7-9.
+- Score 10: The image is explicitly referenced with complete accuracy and proper placement.
+- Example 1: The image is explicitly mentioned and thoroughly explained in the text, with a discussion of NFL play calls perfectly aligned with the visual content.
+The placement and explanation of the image are spot-on, so it warrants a score of 10.
+- Example 2: The image shows a detailed diagram of the human circulatory system, and the text provides an in-depth explanation of how blood flows through the heart and veins.
+The image and text work in perfect harmony to enhance the reader’s understanding.
+The score should be 10.
+- Example 3: The image depicts a vintage car and is referenced explicitly in the context, which discusses the history and evolution of automobile designs.
+The reference is fully accurate, and the image provides essential clarity to the reader, justifying a perfect score of 10.
+Note: The image only needs to be referenced or explained by its context above OR below, not necessarily both.
+Be rigorous and objective in your assessment.
 # Output Instructions
@@
 Figure 20: Prompt template for the image reference evaluation.
-# Task Description
-You are a multi-modal document evaluation assistant.
-You will receive a single image and a multi-modal document written in markdown, where detailed captions serve as proxies for images.
-Your task is to evaluate whether the image should be inserted into the document to enhance understanding for human readers.
-# Images in the Document
-The images in the document have been replaced with their detailed captions. Each image is represented by a single line starting with ‘<image>:’.
-# Document
-markdown
-{text} ‘‘‘ # Image
-<IMAGE_PLACEHOLDER>
-# Scoring Criteria
-Evaluate whether the provided image should be inserted into the document based on the following:
-- If the image is irrelevant to the document’s content or does not aid in comprehension, output a single integer 0, indicating the image should not be inserted.
-- If the image content is too similar to or repetitive of images already represented in the document, output a single integer 0, indicating the image should not be inserted.
-- Otherwise, output a single integer 1, indicating the image should be inserted.
-# Output Instructions
-Your output should contain only two lines:
-1. A brief explanation justifying the score of 0 or 1. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
-Figure 22: Prompt template for the image recall evaluation.
-Appendix E Experimental Setup
+E.2 Evaluation Costs
+The average costs of our evaluation metrics are listed in Table 6.
+Metric Cost (\$ per 1K samples)
+Text-modal Metrics
+Fluency 3.50
+Relevance 3.58
+Context Precision 4.61
+Faithfulness 3.98
+Multi-modal Metrics
+Image Coherence 4.76
+Image Helpfulness 4.92
+Image Reference 4.70
+Table 6: Evaluation costs of our metrics.
+The prices represent the average evaluation cost of the experiments using the joint modeling approach, as presented in Table  2.
+The text-modal metrics correspond to the evaluation cost of an entire generated response, whereas the multi-modal metrics correspond to the evaluation cost per image in the output.
+The Context Precision and Faithfulness metrics are based on the commercial model GPT-4o-mini, while all other metrics use GPT-4o.
+Appendix F Experimental Setup
 Evaluated Models
@@
 GPT-4o GPT-4o
+DeepSeek-V3 Step-1o
 Llama-3.1-70B-Instruct Llama-3.2-90B-Vision-Instruct
@@
 (2023) by using the GPT-4o-mini model.
-Appendix F Case Study
+Appendix G Case Study
 In this section, we provide some cases to demonstrate more details in the M2RAG task:
@@
 Figure 23: A case of comparison between separated and jointly modeling approaches with GPT-4o.
+Appendix H Implementation Details of Model Distillation
+We fine-tune the open-source LLMs Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct and MLLM Qwen2-VL-7B-Instruct to serve as the distilled models.
+To ensure the fine-tuned model effectively captures the comprehensive information embedded in the training corpus, we set the context length to 32,768 tokens during fine-tuning, accommodating the majority of samples within the dataset.
+To optimize the computational efficiency and uphold the performance of the fine-tuned model, we employed Low-Rank Adaptation (LoRA) Hu et al.
+(2021) with the rank of 128 and $\alpha$ of 256.
+Apart from that, we adopt various methods to accelerate training including ZeRO Rajbhandari et al.
+(2020), Flash Attention 2 Dao (2024) and Liger Kernel Hsu et al.
+(2024).
+The model training was conducted on 2 Nvidia A100-SXM4-80GB GPUs with a global batch size of 64 over 3 epochs, resulting in a total of 75 training steps.
+All models are fine-tuned with LLaMA-Factory framework Zheng et al.
+(2024).
+Appendix I Results of Representative Models on the Entire Benchmark Dataset
+The evaluation results of several representative models on the entire benchmark dataset are listed in Table 8.
+Model Type Generation Strategy Model Text-modal Metrics Multi-modal Metrics Overall
+Flu. Rel. CP. Faith. Coher. Help. Ref. Recall
+Reasoner Single OpenAI o3-mini 83.5 90.8 86.8 87.4 74.0 64.6 68.3 84.8 80.0
+Multi OpenAI o3-mini 81.6 84.4 91.1 81.4 69.6 61.1 77.2 100.0 80.8
+LLMs Single GPT-4o 83.7 88.5 90.1 86.5 68.1 57.8 45.0 86.7 75.8
+DeepSeek-V3 83.6 89.0 86.1 84.7 64.4 53.3 44.6 89.6 74.4
+Multi GPT-4o 82.8 86.0 91.4 82.9 69.6 61.1 77.4 100.0 81.4
+DeepSeek-V3 82.4 85.8 90.6 84.8 67.9 59.5 76.2 99.8 80.9
+MLLMs Single GPT-4o 84.1 88.9 89.4 85.6 69.1 58.2 44.6 87.3 75.9
+Step-1o 88.0 91.1 89.9 86.5 65.5 54.1 47.7 89.0 76.5
+Multi GPT-4o 82.4 86.8 91.0 81.7 68.5 59.4 76.8 100.0 80.8
+Step-1o 84.0 88.1 92.2 86.7 69.0 60.4 75.9 97.7 81.7
+Fine-tuned Models
+LLMs Single Llama-3.1-8B-Instruct 76.1 80.5 92.6 79.8 67.0 59.6 76.0 98.4 78.7
+Qwen2.5-7B-Instruct 78.0 81.1 89.0 79.6 66.6 58.8 75.3 93.4 77.7
+MLLM Single Qwen2-VL-7B-Instruct 61.7 77.2 87.2 68.4 64.3 57.2 72.6 95.4 73.0
+Table 8: The overall experiment results of representative models on the entire benchmark dataset.